### v1

# Brauer groups of tame stacky curves and $\mu_{r}$-gerbes over them

Abstract.
We fit the Brauer group of a $\mu_{r}$-gerbe over a (possibly arbitrarily singular) stacky curve into an exact sequence and give characterizations for when it is short exact and a condition for when it splits. We also give a precise formula for the Brauer group of a $\mu_{r}$-gerbe in the case that the base curve is smooth. We end by showing that $\operatorname{Br}=\operatorname{Br}^{\prime}=\operatorname{H}^{2}$ for separated tame one-dimensional algebraic stacks of finite type over a field.
1. Introduction
Brauer groups have been the object of intense study since they were first defined by Brauer in the 1920’s. Brauer groups have a near endless amount of applications, and in fact this paper had its origins as a computational tool used in an upcoming paper with William Newman.
The fundamental aim of this paper is to provide tools and techniques with which to compute Brauer groups of one-dimensional Deligne-Mumford stacks over algebraically closed fields (in fact, the techniques work in more generality than this: see the final paragraph before Section 1.1). To this end, our primary results are: complete computations of the Brauer groups of stacky curves; complete computations of the Brauer groups of $\mu_{r}$-gerbes over smooth stacky curves; and an exact sequence for the Brauer group of a $\mu_{r}$-gerbe over a stacky curve along with precise conditions for when it is short exact.
The reader who has looked ahead to page 2 will notice that, though we say Brauer group, all our results are stated for $\operatorname{H}^{2}(-,\mathbb{G}_{m})$. This is because this paper also provides a new class of stacks for which $\operatorname{Br}=\operatorname{Br}^{\prime}=\operatorname{H}^{2}$: separated tame one-dimensional algebraic stacks of finite type over a field.
The main inspiration for, and the fundamental technical basis of, this work is the results in the note [Mei18]. Inspiration was also drawn from [AM20, Lie11, Shi21]. While preparing this paper, the author learned of the work [Ach24] which uses a similar approach (leveraging the results of Section 3.1 in conjunction with spectral sequences), although the papers operate on different levels of generality. The paper [Ach24] makes weaker assumptions on the base stack and the étale-ness of stabilizers, offering techniques which work for a wide variety of stacks provided they are locally Brauerless ([Ach24, Definition 4.13]), while this paper works in the more restrictive setting of one-dimensional DM stacks over algebraically closed fields, but with these stronger assumptions is able to state results for stacks which are highly singular and not locally Brauerless. In short, we insist that our stabilizers must be étale, but impose no other conditions. Whereas being locally Brauerless allows for non-étale stabilizers but places additional requirements on the étale part. In this way, these two papers work in complementary levels of generality. The research in this paper, outside of Section 5, was conducted independently of [Ach24].
1.1. Main results
Our main result is the following thorough classification of Brauer groups of $\mu_{r}$-gerbes over stacky curves. Note that here $\mathcal{C}$ is allowed to be arbitrarily singular.
Theorem (Theorem 4.15).
Suppose that $\mathcal{C}$ is a tame separated one-dimensional Deligne-Mumford stack with trivial generic stabilizer of finite type over an algebraically closed field $\mathbbm{k}$. Let $f:\mathcal{C}\rightarrow C$ be the coarse moduli space of $\mathcal{C}$. Suppose that $p_{1},\dots,p_{n}$ are the geometric points in $C$ whose preimage in $\mathcal{C}$ have non-trivial stabilizers $G_{1},\dots,G_{n}$. Let $\pi:\mathcal{G}\rightarrow\mathcal{C}$ be a $\mu_{r}$-gerbe, and denote the stabilizers in $\mathcal{G}$ over $p_{1},\dots,p_{n}$ as $E_{1},\dots,E_{n}$. Then the following sequence is exact
$\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r).$
Moreover,
  1. (1)
    the left map is injective if and only if $\mathcal{G}$ is a root gerbe over $\mathcal{C}$,
  2. (2)
    $\mathcal{G}$ is a root gerbe over $\mathcal{C}$ if and only if it is a root gerbe on each fiber $BE_{i}\rightarrow BG_{i}$,
  3. (3)
    the right map is surjective if and only if
    $\inf:\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$
    is injective for all $i=1,\dots,n$, and
  4. (4)
    if $\gcd(r,\left|G_{i}\right|)=1$ for all $i=1,\dots,n$, then:
    1. (a)
      $\pi:\mathcal{G}\rightarrow\mathcal{C}$ is automatically a root gerbe,
    2. (b)
      the inflation map is automatically injective, and
    3. (c)
      we have the following splitting
      $\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})=\operatorname{H}^{1}(C,\mathbb{Z}/r)\oplus\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times}).$
In the case where $\mathcal{C}$ is smooth, this vastly simplifies into the following (the interested reader can find more cases in Section 4.4).
Theorem (Corollary 4.16).
Suppose that $\mathcal{C}$ is a smooth separated tame one-dimensional Deligne-Mumford stack with trivial generic stabilizer of finite type over an algebraically closed field $\mathbbm{k}$. Let $\pi:\mathcal{G}\rightarrow\mathcal{C}$ be a $\mu_{r}$-gerbe. Then
$\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})=\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r).$
We also present a comprehensive computation of the cohomology of $\mathbb{G}_{m}$ on $\mathcal{C}$, with no smoothness requirements.
Theorem (Proposition 4.5).
Let $\mathcal{C}$ be a separated tame one-dimensional Deligne-Mumford stack with trivial generic stabilizer of finite type over an algebraically closed field $\mathbbm{k}$. Let $G_{1},\dots,G_{n}$ be the non-trivial stabilizers of geometric points of $\mathcal{C}$. Then for all $k\geq 2$,
$\operatorname{H}^{k}(\mathcal{C},\mathbb{G}_{m})=\bigoplus_{i=1}^{n}\operatorname{H}^{k}(G_{i},\mathbbm{k}^{\times}).$
At the very end of the paper, we apply our methods to extend the class of stacks for which $\operatorname{Br}=\operatorname{Br}^{\prime}=\operatorname{H}^{2}$ is known.
Theorem (Theorem 3.4).
Let $\mathcal{X}$ be a separated tame one-dimensional algebraic stack of finite type over a field $\mathbbm{k}$ (here we do not assume that $\mathbbm{k}$ is algebraically closed). Then
$\operatorname{Br}(\mathcal{X})=\operatorname{Br}^{\prime}(\mathcal{X})=\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m}).$
Along the way, we also collect multiple characterizations of when a $\mu_{r}$-gerbe over a stacky curve is a root gerbe (see Proposition 3.8 and Corollaries 4.8 and 4.9).
1.2. Conventions
We work over a fixed algebraically closed field $\mathbbm{k}$, except in Section 5, where $\mathbbm{k}$ is not assumed to be algebraically closed. All stacks are assumed to be tame, separated, Deligne-Mumford, and of finite type over $\mathbbm{k}$, except for in Section 5, where we replace Deligne-Mumford with algebraic. When $\mathcal{X}$ is one-dimensional, we will refer to $\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})$ as the Brauer group of $\mathcal{X}$ (see Note 3.5). A Deligne-Mumford stack is tame if $\operatorname{char}\mathbbm{k}$ does not divide the order of any geometric stabilizer group, and an algebraic stack is tame if it has finite inertia and all geometric stabilizers are linearly reductive (see [AOV08]).
We take a stacky curve to be a separated one-dimensional Deligne-Mumford stack with trivial generic stabilizer of finite type over a field, though we tend to avoid this terminology (the eagle-eyed reader who has seen the title of this paper will notice that we failed this immediately). Instead, we usually directly state the conditions we are assuming, since there are conflicting definitions of stacky curve in the literature.
1.3. Acknowledgments
This project was born out of an upcoming paper with William Newman on ¯$\overline{\mathcal{M}}_{1,1}(B\mu_{2})$, the moduli stack of twisted stable maps from an elliptic curve to $B\mu_{2}$. I am heavily indebted to him for hours of hard work and insightful comments on draft versions of this paper. The results of this paper will be used in this upcoming work to both help understand the gerbe structure of ¯$\overline{\mathcal{M}}_{1,1}(B\mu_{2})$ as well as to compute the Picard groups of nodal stacky curves.
I would also like to thank Cherry Ng for helpful conversations on group cohomology, Jackson Morris for fantastic advice on how to compute using spectral sequences, and Minseon Shin for suggesting a simplification of an early version of Proposition 3.14.
2. Background
This paper relies very heavily on the careful application of both spectral sequences and group cohomology (which is related by the results of Section 3.1 to étale cohomology). Therefore we consider it appropriate to give a brief explanation of the properties and results we will be invoking.
2.1. Group cohomology
For a group $G$ and an abelian group $A$ on which $G$ acts additively (called a $G$-module), define the invariant subgroup
$A^{G}=\{a\in A:ga=a,\ \forall g\in G\}.$
Then the cohomology of $G$ with coefficients in $A$, $\operatorname{H}^{n}(G,A)$, is the right derived functor $\mathbf{R}^{n}(-^{G})(A)$.
Proposition 2.1.
Let $G$ be a group and $A$ a $G$-module. If $G$ is $m$-torsion, then so is $\operatorname{H}^{i}(G,A)$ for all $i\geq 1$. If $A$ is $n$-torsion, then so is $\operatorname{H}^{i}(G,A)$ for all $i\geq 0$.
Proposition 2.2.
Let $\mathbbm{k}$ be an algebraically closed field. Let $\mu_{r}$ act trivially on $\mathbbm{k}^{\times}$. Then
$\operatorname{H}^{2k-1}(\mu_{r},\mathbbm{k}^{\times})=\operatorname{Hom}(\mu_{r},\mathbbm{k}^{\times})=\mathbb{Z}/r$
and
$\operatorname{H}^{2k}(\mu_{r},\mathbbm{k}^{\times})=0$
for $k\geq 1$.
Proposition 2.3.
Let $\mathbb{Z}/n$ and $\mathbb{Z}/mn$ act trivially on $\mathbbm{k}^{\times}$. Then
$\operatorname{H}^{3}(\mathbb{Z}/n,\mathbbm{k}^{\times})\hookrightarrow\operatorname{H}^{3}(\mathbb{Z}/mn,\mathbbm{k}^{\times}),$
given by the canonical injection $1\mapsto m$.
Proof.
More generally, the map
$\operatorname{H}^{k}(\mathbb{Z}/n,\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{k}(\mathbb{Z}/mn,\mathbbm{k}^{\times})$
is given by $1\mapsto m^{\left\lfloor{\frac{k}{2}}\right\rfloor}$ [Ser79, XIII Ex. 1.1], which for $k=3$ is $1\mapsto m$. ∎
Proposition 2.4.
If $\mathbb{Z}/a$ and $\mathbb{Z}/b$ both act trivially on $\mathbbm{k}^{\times}$, then
$\operatorname{H}^{k}(\mathbb{Z}/a,\mathbbm{k}^{\times})\hookrightarrow\operatorname{H}^{k}(\mathbb{Z}/a\oplus\mathbb{Z}/b,\mathbbm{k}^{\times})$
for all $k\geq 1$.
2.2. Spectral Sequences
A standard reference for this section is [Wei94, Chapter 5].
Definition 2.5.
A spectral sequence consists of the following data
  1. (1)
    sheets $E_{r}$ for $r\geq r_{0}$,
  2. (2)
    for all integers $p,q$, objects $E_{r}^{p,q}$, and
  3. (3)
    maps $d_{r}:E_{r}^{p,q}\rightarrow E_{r}^{p+r,q-(r-1)}$ such that $d_{r}\circ d_{r}=0$ and $E_{r+1}^{p,q}$ is the homology of $E_{r}$ at $(p,q)$.
In practice, many spectral sequences (including the ones used in this paper), begin at the $E_{2}$ sheet. One then generally uses the $E_{2}$ sheet to calculate the subsequent sheets.
In this section we will assume that all spectral sequences are first quadrant, meaning the only non-zero terms have $p,q\geq 0$. Note that any first quadrant spectral sequence must eventually stabilize in the sense that there is some $N$ such that $E_{r}^{p,q}=E_{N}^{p,q}$ for all $r\geq N$. We call this stable value $E_{\infty}^{p,q}$.
We say that a first quadrant spectral sequence converges to $\operatorname{H}^{*}$ for some collection of objects $(\operatorname{H}^{n})$ if for each $n$ there is a filtration
$0\subseteq F^{n}\operatorname{H}^{n}\subseteq F^{n-1}\operatorname{H}^{n}\subseteq\dots\subseteq F^{0}\operatorname{H}^{n}=\operatorname{H}^{n}$
such that
$E_{\infty}^{p,q}\cong F^{p}\operatorname{H}^{p+q}/F^{p+1}\operatorname{H}^{p+q}.$
When such a filtration exists, we denote this by
$E_{r}^{p,q}\implies\operatorname{H}^{p+q}.$
Convergent first quadrant spectral sequences always have a sequence of low-degree terms.
Lemma 2.6 (Sequence of low-degree terms).
If
$E_{2}^{p,q}\implies\operatorname{H}^{p+q}(X)$
is a first quadrant spectral sequence, then there is an exact sequence
$0\rightarrow E_{2}^{1,0}\rightarrow\operatorname{H}^{1}(X)\rightarrow E_{2}^{0,1}\rightarrow E_{2}^{2,0}\rightarrow\ker\left[\operatorname{H}^{2}(X)\rightarrow E_{2}^{0,2}\right]\rightarrow E_{2}^{1,1}\rightarrow E_{2}^{3,0}.$
This can be extended an additional term to $\operatorname{H}^{3}(X)$ if $E_{2}^{0,2}=0$.
Proof.
This is standard, but we will give an explanation of the extension to $\operatorname{H}^{3}(X)$. The filtration on $\operatorname{H}^{3}(X)$ associated to the spectral sequence is
$0\subseteq F^{3}\operatorname{H}^{3}\subseteq F^{2}\operatorname{H}^{3}\subseteq F^{1}\operatorname{H}^{3}\subseteq F^{0}\operatorname{H}^{3}=\operatorname{H}^{3}(X),$
where $F^{i}\operatorname{H}^{3}/F^{i+1}\operatorname{H}^{3}\cong E_{\infty}^{i,3-i}$. Therefore there is a canonical injection $E_{\infty}^{3,0}\hookrightarrow\operatorname{H}^{3}(X)$. Exactness of our desired extension says that we need a morphism $E_{2}^{3,0}\rightarrow\operatorname{H}^{3}(X)$ whose kernel is precisely the image of $E_{2}^{1,1}$.
Note that
$E_{3}^{3,0}=\frac{\ker(E_{2}^{3,0}\rightarrow E_{2}^{5,-1}=0)}{\operatorname{im}(E_{2}^{1,1}\rightarrow E_{2}^{3,0})}$
and so we have exactness of the extension as long as $E_{3}^{3,0}=E_{\infty}^{3,0}$. The obstruction to this happening is whether or not $E_{4}^{3,0}=E_{3}^{3,0}$. Since
$E_{4}^{3,0}=\frac{\ker(E_{3}^{3,0}\rightarrow E_{3}^{6,-2}=0)}{\operatorname{im}(E_{3}^{0,2}\rightarrow E_{3}^{3,0})},$
we see that $E_{4}^{3,0}=E_{3}^{3,0}$ if and only if the image of $E_{3}^{0,2}$ is trivial in $E_{3}^{3,0}$. In particular, we get equality and hence a further extension of the sequence if $E_{2}^{0,2}=0$. ∎
We will make extensive use of the Grothendieck spectral sequence, the Leray spectral sequence, and the Lyndon-Hochschild-Serre spectral sequence. Our statement of the Grothendieck spectral sequence is only for the context in which we will use it (pushforward of abelian sheaves on stacks), though it applies in much larger generality.
Definition 2.7.
For morphisms $f:\mathcal{X}\rightarrow\mathcal{Y}$ and $g:\mathcal{Y}\rightarrow\mathcal{Z}$ of stacks and an abelian sheaf $F$ on $\mathcal{X}$, the Grothendieck spectral sequence is
$E_{2}^{p,q}=\mathbf{R}^{p}g_{*}(\mathbf{R}^{q}f_{*}F)\implies\mathbf{R}^{p+q}(g\circ f)_{*}F.$
Definition 2.8.
For a morphism $f:\mathcal{X}\rightarrow\mathcal{Y}$ of stacks and an abelian sheaf $F$ on $\mathcal{X}$, the Leray spectral sequence is
$E_{2}^{p,q}=\operatorname{H}^{p}(\mathcal{Y},\mathbf{R}^{q}f_{*}F)\implies\operatorname{H}^{p+q}(\mathcal{X},F).$
Definition 2.9.
For a group $G$ with normal subgroup $N$ and $G$-module $A$, the Lyndon-Hochschild-Serre spectral sequence is
$E_{2}^{p,q}=\operatorname{H}^{p}(G/N,\operatorname{H}^{q}(N,A))\implies\operatorname{H}^{p+q}(G,A).$
3. Basics on gerbes
Definition 3.1.
A stack $\mathcal{X}$ over a site $\mathcal{S}$ is called a gerbe if
  1. (1)
    for every object $T\in\mathcal{S}$, there exists a covering $\{T_{i}\rightarrow T\}$ in $\mathcal{S}$ such that each $\mathcal{X}(T_{i})$ is non-empty, and
  2. (2)
    for objects $x,y\in\mathcal{X}$ over $T\in\mathcal{S}$, there exists a covering $\{T_{i}\rightarrow T\}$ and isomorphisms $x|_{T_{i}}\rightarrow y|_{T_{i}}$ for each $i$.
Definition 3.2.
Let $G$ be an abelian sheaf on $\mathcal{S}$. A stack $\mathcal{X}$ over $\mathcal{S}$ is called a $G$-gerbe (or a banded $G$-gerbe) if it is a gerbe together with the data of isomorphisms $\psi_{x}:G|_{T}\rightarrow\operatorname{Aut}_{T}(x)$ of sheaves for each object $x\in\mathcal{X}(T)$. We require that for each isomorphism $\alpha:x\xrightarrow{\sim}y$ over $T$, the diagram
[…]
commutes.
Loosely speaking, this says that $G$-gerbes are a generalization of the notion of classifying stack, in the sense that a $G$-gerbe is a stack which locally is isomorphic to $BG$, much in the same way that $G$-torsors are locally isomorphic to $G$. It is a foundational fact that $G$-gerbes are classified by $\operatorname{H}^{2}(\mathcal{S},G)$ ([Ols16, Theorem 12.2.8]).
Definition 3.3.
We define the cohomological Brauer group of a stack $\mathcal{X}$ to be $\operatorname{Br}^{\prime}(\mathcal{X}):=\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})_{tors}$.
Theorem 3.4.
Let $\mathcal{X}$ be a separated tame one-dimensional algebraic stack of finite type over a field $\mathbbm{k}$ (here we do not assume that $\mathbbm{k}$ is algebraically closed). Then
$\operatorname{Br}(\mathcal{X})=\operatorname{Br}^{\prime}(\mathcal{X})=\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m}).$
Proof.
See Section 5. ∎
Note 3.5.
As stated in Section 1.2, in light of this theorem we will refer to $\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})$ as the Brauer group of $\mathcal{X}$.
Proposition 3.6.
Let $\pi:\mathcal{G}\rightarrow X$ be a $\mu_{r}$-gerbe over a Deligne-Mumford stack. Then there is a natural isomorphism
$\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m}\cong\operatorname{Hom}(\mu_{r},\mathbb{G}_{m})=\mathbb{Z}/r.$
Proof.
Observe that to each line bundle $L\in\operatorname{Pic}(\mathcal{G})$ we may associate a character $\chi_{L}\in\operatorname{Hom}(\mu_{r},\mathbb{G}_{m})$. For each étale $V\rightarrow X$, we have the exact sequence
$0\rightarrow\operatorname{Pic}(V)\rightarrow\operatorname{Pic}(\mathcal{G}|_{V})\xrightarrow{L\mapsto\chi_{L}}\operatorname{Hom}(\mu_{r},\mathbb{G}_{m}).$
Since $\mathcal{G}$ is étale locally trivial, this sequence is étale locally surjective on the right. That is, $\operatorname{Hom}(\mu_{r},\mathbb{G}_{m})$ is precisely the sheafification of $\operatorname{Pic}(\mathcal{G})/\operatorname{Pic}(X)$, which is $\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m}$. ∎
Our principal gerbes of study will be root gerbes, which were introduced in the literature in [Cad07].
Definition 3.7.
Let $X$ be an algebraic stack. Then the data of a line bundle $L$ on $X$ is equivalent to the data of a morphism $X\xrightarrow{[L]}B\mathbb{G}_{m}$. Letting $r$ denote the $r^{\text{th}}$ power map $r:\mathbb{G}_{m}\rightarrow\mathbb{G}_{m}$, the $r^{\text{th}}$ root gerbe, written $X(\sqrt[r]{L})$, is defined as the following fiber product
[…]
A morphism from a scheme $T$ to $X(\sqrt[r]{L})$ is equivalent to the data of a morphism $f:T\rightarrow X$, a line bundle $L^{\prime}$ on $T$, and an isomorphism $L^{\prime\otimes n}\cong f^{*}(L)$.
Proposition 3.8.
Let $X$ be a Deligne-Mumford stack. Then a $\mu_{r}$-gerbe $\pi:\mathcal{X}\rightarrow X$ is a root gerbe if and only if $\operatorname{H}^{2}(X,\mathbb{G}_{m})\rightarrow\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})$ is injective.
Proof.
Consider the Leray spectral sequence
$E_{2}^{p,q}=\operatorname{H}^{p}(X,\mathbf{R}^{q}\pi_{*}\mathbb{G}_{m})\implies\operatorname{H}^{p+q}(\mathcal{X},\mathbb{G}_{m}).$
Since $\pi_{*}\mathbb{G}_{m}=\mathbb{G}_{m}$, the beginning of the associated sequence of low-degree terms is
$0\rightarrow\operatorname{H}^{1}(X,\mathbb{G}_{m})\rightarrow\operatorname{H}^{1}(\mathcal{X},\mathbb{G}_{m})\rightarrow\operatorname{H}^{0}(X,\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m})=$
$\operatorname{H}^{0}(X,\operatorname{Hom}(\mu_{r},\mathbb{G}_{m}))\xrightarrow{d_{2}}\operatorname{H}^{2}(X,\mathbb{G}_{m})\xrightarrow{\pi^{*}}\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m}).$
Therefore the injectivity of $\pi^{*}$ is equivalent to the triviality of $d_{2}$. By [Lop23, Theorem 1.2], $d_{2}$ is just the map which sends $\epsilon:\mu_{r}\rightarrow\mathbb{G}_{m}$ to the image of $[\mathcal{G}]\in\operatorname{H}^{2}(X,\mu_{r})$ under the pushforward map
$\epsilon_{*}:\operatorname{H}^{2}(X,\mu_{r})\rightarrow\operatorname{H}^{2}(X,\mathbb{G}_{m}).$
Setting $\epsilon$ to be the canonical inclusion $\mu_{r}\hookrightarrow\mathbb{G}_{m}$, the Kummer sequence shows that $d_{2}$ is trivial if and only if $[\mathcal{G}]$ is in the image of $\operatorname{H}^{1}(X,\mathbb{G}_{m})$. That is, if and only if $\mathcal{G}$ is a root gerbe. ∎
Proposition 3.9 ([AM20], Proposition 2.5 (iv)).
If $X$ is a regular and noetherian Deligne-Mumford stack, and if $U\subseteq X$ is a dense open subset, then the restriction map $\operatorname{H}^{2}(X,\mathbb{G}_{m})\rightarrow\operatorname{H}^{2}(U,\mathbb{G}_{m})$ is injective.
Theorem 3.10 (Tsen’s Theorem).
Let $X$ be a one-dimensional scheme over an algebraically closed field $\mathbbm{k}$. Then $\operatorname{H}^{i}(X,\mathbb{G}_{m})=0$ for $i\geq 2$.
Theorem 3.11 (Stacky Tsen’s Theorem).
Let $\mathcal{C}$ be a smooth separated Deligne-Mumford stack with trivial generic stabilizer of finite type over an algebraically closed field. Then $\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})=0$.
Proof.
Since there are only finitely many points which have non-trivial stabilizer, this follows from Tsen’s Theorem and Proposition 3.9. ∎
Note 3.12.
Note the contrast between Theorem 3.10, which applies to all curves, and Theorem 3.11, which applies only when $\mathcal{C}$ is smooth. Intuitively, this says that for a stacky curve (which is generically schemey), any non-triviality in $\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})$ must come from stabilizers at singularities, which is precisely what we will see in Proposition 4.5 and Note 4.6.
3.1. Some results of Meier
The following results, with small adjustments to fit our context, appear in [ACV03] and [Mei18].
Definition 3.13 ([ATWo20], 2.3.1).
A morphism $f:X\rightarrow Y$ of stacks is called a coarsening if the inertia $\mathcal{I}_{X/Y}$ is finite over $X$ and for any flat $Z\rightarrow Y$ from an algebraic space $Z$ the resulting base change $X\times_{Y}Z\rightarrow Z$ is a coarse moduli space.
Proposition 3.14 ([ACV03], Proposition A.0.1 [Mei18], Proposition 8).
Let $\mathcal{X}$ be a separated tame Deligne-Mumford stack of finite type over $\mathbbm{k}$ with coarsening $f:\mathcal{X}\rightarrow X$. Let $x:\operatorname{Spec}K\rightarrow X$ be a geometric point. Let $G=(\mathcal{I}_{\mathcal{X}/X})_{x}$ be the relative stabilizer. Then for all $p\geq 0$
$(\mathbf{R}^{p}f_{*}\mathbb{G}_{m})_{x}=\operatorname{H}^{p}(G,R^{\times})$
for some strictly Henselian local ring $R$ with residue field $K$. Moreover, the group $G$ acts trivially on $K$.
Proof.
This is originally stated for coarse moduli spaces. We reduce to that case by observing that since $\mathcal{X}\rightarrow X$ is a coarsening we may restrict to any étale neighborhood from a scheme $V\rightarrow X$ and then be in the setting of a coarse moduli space $\mathcal{X}_{V}\rightarrow V$.
[ACV03, Proposition A.0.1] provides an equality
$(\mathbf{R}^{p}f_{*}\mathbb{G}_{m})_{x}=\operatorname{H}^{p}(G,(\mathbb{G}_{m})_{x^{\prime}}),$
where $x^{\prime}$ is the preimage of $x$ in $\mathcal{X}$, and [Mei18, Proposition 8] shows it is of the form $\operatorname{H}^{p}(G,R^{\times})$ with trivial action on $K$. ∎
Corollary 3.15 ([Mei18], Lemma 9).
Let $\mathcal{X}$ and $G$ be as in Proposition 3.14. Then for all geometric points $x:\operatorname{Spec}K\rightarrow X$ and for all $n\geq 2$,
$(\mathbf{R}^{n}f_{*}\mathbb{G}_{m})_{x}=\operatorname{H}^{n}(G,K^{\times}).$
Note 3.16.
Meier states this lemma without direct reference to stalks, though it is implicit. It is also only stated for the case $n=2$ and only for the $\ell$-torsion with $\ell\neq\operatorname{char}K$. The restrictions on $n$ and the $\ell$-torsion originate from trying to show that $\operatorname{H}^{n}(G,R^{\times})=\operatorname{H}^{n}(G,K^{\times})$. However, the proof actually shows that the $\ell$-torsion agrees for all $n\geq 2$, and the issue with $(\operatorname{char}K)$-torsion disappears once we require that $|G|$ and $\operatorname{char}K$ are coprime, since $\operatorname{H}^{n}(G,-)$ is $|G|$-torsion for all $n\geq 1$. But $|G|$ and $\operatorname{char}K$ are coprime when $\mathcal{X}$ is tame and Deligne-Mumford, which is precisely our setting.
We add here an additional result inspired by the above.
Proposition 3.17.
Let $\mathcal{X}$ and $G$ be as in Proposition 3.14. Then for all geometric points $x:\operatorname{Spec}K\rightarrow X$ and all $n\geq 1$,
$(\mathbf{R}^{n}f_{*}\mu_{r})_{x}=\operatorname{H}^{n}(G,\mu_{r}(K)).$
Proof.
This reduces to showing that if $\mathcal{X}=[\operatorname{Spec}R/G]$, then $\operatorname{H}^{n}(\mathcal{X},\mu_{r})=\operatorname{H}^{n}(G,\mu_{r})$. By the descent spectral sequence we have
$\operatorname{H}^{n}(\mathcal{X},\mu_{r})=\operatorname{H}^{n}(G,\mu_{r}(R))=\operatorname{H}^{n}(G,\mu_{r}(K)),$
where the last equality comes from $R$ being a strictly Henselian local ring. ∎
4. The Brauer group of a $\mu_{r}$-gerbe
In this section, we fix the following data:
  • $\mathcal{C}$ is a tame one-dimensional separated Deligne-Mumford stack with trivial generic stabilizer of finite type over an algebraically closed field $\mathbbm{k}$,
  • $\pi:\mathcal{G}\rightarrow\mathcal{C}$ is a $\mu_{r}$-gerbe,
  • $g:\mathcal{G}\rightarrow C$ is the coarse moduli space (note that $C$ is the coarse moduli space of both $\mathcal{G}$ and $\mathcal{C}$),
  • $p_{1},\dots,p_{n}$ are the points of $C$ whose preimage in $\mathcal{C}$ have non-trivial stabilizers $G_{i}$, and
  • the preimages of $p_{1},\dots,p_{n}$ in $\mathcal{G}$ have stabilizers $E_{1},\dots,E_{n}$, so that for each $i$ we have the exact sequence of groups
    $0\rightarrow\mu_{r}\rightarrow E_{i}\rightarrow G_{i}\rightarrow 0$
This can be summarized in the diagram
${\mathcal{G}}$${\mathcal{C}}$${C.}$$\scriptstyle{\pi}$$\scriptstyle{g}$$\scriptstyle{f}$
Consider the Leray spectral sequence for the sheaf $\mathbb{G}_{m,\mathcal{G}}$ and the map $\pi:\mathcal{G}\rightarrow\mathcal{C}$:
$E_{2}^{p,q}=\operatorname{H}^{p}(\mathcal{C},\mathbf{R}^{q}\pi_{*}\mathbb{G}_{m,\mathcal{G}})\implies\operatorname{H}^{p+q}(\mathcal{G},\mathbb{G}_{m,\mathcal{G}}).$
Using the fact that $\pi_{*}\mathbb{G}_{m,\mathcal{G}}=\mathbb{G}_{m,\mathcal{C}}$, the associated sequence of low-degree terms is
(1) $0\rightarrow\operatorname{H}^{1}(\mathcal{C},\mathbb{G}_{m})\rightarrow\operatorname{H}^{1}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{0}(\mathcal{C},\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m})\rightarrow\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})\rightarrow\\ \ker\left[\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{0}(\mathcal{C},\mathbf{R}^{2}\pi_{*}\mathbb{G}_{m})\right]\rightarrow\operatorname{H}^{1}(\mathcal{C},\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m})\rightarrow\operatorname{H}^{3}(\mathcal{C},\mathbb{G}_{m}).$
In this section we will thoroughly analyze Sequence 1 to extract a short exact sequence containing $\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})$ and then give conditions under which it splits and simplifies. We open by using the Grothendieck spectral sequence to relate $\mathbf{R}^{*}f_{*}\mathbb{G}_{m}$, $\mathbf{R}^{*}g_{*}\mathbb{G}_{m}$, and $\mathbf{R}^{*}\pi_{*}\mathbb{G}_{m}$.
4.1. Grothendieck spectral sequence
To understand $\mathbf{R}^{*}g_{*}\mathbb{G}_{m}$, we consider the Grothendieck spectral sequence for the composition $g=f\circ\pi$ and the sheaf $\mathbb{G}_{m}$
$E_{2}^{pq}=\left(\mathbf{R}^{p}f_{*}\circ\mathbf{R}^{q}\pi_{*}\right)(\mathbb{G}_{m})\implies\mathbf{R}^{p+q}g_{*}(\mathbb{G}_{m}).$
Accounting for the fact that the pushforward of $\mathbb{G}_{m}$ is $\mathbb{G}_{m}$, the beginning of the sequence of low-degree terms is
[…]
We may extract two short exact sequences from this:
$0\rightarrow\mathbf{R}^{1}f_{*}\mathbb{G}_{m}\rightarrow\mathbf{R}^{1}g_{*}\mathbb{G}_{m}\rightarrow\ker\gamma\rightarrow 0$
and
$0\rightarrow\ker\gamma\rightarrow f_{*}(\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m})\rightarrow\ker\delta\rightarrow 0$
Lemma 4.1.
For all $i\geq 2$,
$\operatorname{H}^{i}(C,\mathbf{R}^{1}g_{*}\mathbb{G}_{m})\cong\operatorname{H}^{i}(C,f_{*}(\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m}))$
and the isomorphism holds for $i=1$ if and only if
$\operatorname{H}^{0}(C,f_{*}(\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m}))\rightarrow\operatorname{H}^{0}(C,\ker\delta)$
is surjective.
Proof.
Notice that since $\mathcal{C}$ has trivial generic stabilizer, Corollary 3.15 says that for $i\geq 1$ the sheaves $\mathbf{R}^{i}f_{*}\mathbb{G}_{m}$ are supported on a finite set (the points where $\mathcal{C}$ has non-trivial stack structure), and so their higher cohomology vanishes. Hence for $i\geq 1$
$0\rightarrow\operatorname{H}^{i}(C,\mathbf{R}^{1}g_{*}\mathbb{G}_{m})\rightarrow\operatorname{H}^{i}(C,\ker\gamma)\rightarrow 0$
That is,
$\operatorname{H}^{i}(C,\mathbf{R}^{1}g_{*}\mathbb{G}_{m})\cong\operatorname{H}^{i}(C,\ker\gamma).$
But using the second exact sequence from above gives, in degree 0 and 1,
$\operatorname{H}^{0}(C,f_{*}(\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m}))\rightarrow\operatorname{H}^{0}(C,\ker\delta)\rightarrow\operatorname{H}^{1}(C,\ker\gamma)\rightarrow\operatorname{H}^{1}(C,f_{*}(\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m}))\rightarrow 0$
and, in degree above 1,
$0\rightarrow\operatorname{H}^{i}(C,\ker\gamma)\rightarrow\operatorname{H}^{i}(C,f_{*}(\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m}))\rightarrow 0$
This establishes that for all $i\geq 2$ we have
$\operatorname{H}^{i}(C,\mathbf{R}^{1}g_{*}\mathbb{G}_{m})\cong\operatorname{H}^{i}(C,\ker\gamma)\cong\operatorname{H}^{i}(C,f_{*}(\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m}))$
and that the isomorphism holds for $i=1$ if and only if
$\operatorname{H}^{0}(C,f_{*}(\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m}))\rightarrow\operatorname{H}^{0}(C,\ker\delta)$
is surjective. ∎
Note 4.2.
In the case where
$\mathbf{R}^{2}f_{*}\mathbb{G}_{m}\rightarrow\mathbf{R}^{2}g_{*}\mathbb{G}_{m}$
is injective, the first statement of Lemma 4.1 holds for all $i\geq 1$. We will see in Corollary 4.9 that this happens precisely when $\mathcal{G}$ is a root gerbe over $\mathcal{C}$.
Lemma 4.3.
For all $k\geq 1$,
$\mathbf{R}^{2k}g_{*}\mathbb{G}_{m}=\bigoplus_{i=1}^{n}p_{i_{*}}\operatorname{H}^{2k}(E_{i},\mathbbm{k}^{\times}).$
Proof.
By Corollary 3.15 and Proposition 2.2, for all $x\neq p_{1},\dots,p_{n}$ we have $(\mathbf{R}^{2k}g_{*}\mathbb{G}_{m})_{x}=\operatorname{H}^{2k}(\mu_{r},\mathbbm{k}^{\times})=0$. Therefore $\mathbf{R}^{2k}g_{*}\mathbb{G}_{m}$ is supported at the $p_{i}$’s and hence
$\mathbf{R}^{2k}g_{*}\mathbb{G}_{m}=\bigoplus_{i=1}^{n}(\mathbf{R}^{2k}g_{*}\mathbb{G}_{m})_{p_{i}}=\bigoplus_{i=1}^{n}p_{i_{*}}\operatorname{H}^{2k}(E_{i},\mathbbm{k}^{\times}),$
by Corollary 3.15. ∎
4.2. The cohomology of $\mathcal{C}$
Lemma 4.4.
For all $k\geq 2$,
$\mathbf{R}^{k}f_{*}\mathbb{G}_{m}=\bigoplus_{i=1}^{n}p_{i_{*}}\operatorname{H}^{k}(G_{i},\mathbbm{k}^{\times}).$
Proof.
Let $x$ be a geometric point of $C$ and $G_{x}$ the stabilizer of the preimage of $x$ in $\mathcal{C}$. By Corollary 3.15, $(\mathbf{R}^{k}f_{*}\mathbb{G}_{m})_{x}=\operatorname{H}^{k}(G_{x},\mathbbm{k}^{\times})$, and this vanishes away from $p_{1},\dots,p_{n}$ since $G_{x}=0$ generically. Hence $\mathbf{R}^{k}f_{*}\mathbb{G}_{m}$ is supported at $p_{1},\dots,p_{n}$, and the lemma follows. ∎
Proposition 4.5.
For all $k\geq 2$ we have that
$\operatorname{H}^{k}(\mathcal{C},\mathbb{G}_{m})=\operatorname{H}^{0}(C,\mathbf{R}^{k}f_{*}\mathbb{G}_{m})=\bigoplus_{i=1}^{n}\operatorname{H}^{k}(G_{i},\mathbbm{k}^{\times}).$
Proof.
We use the Leray spectral sequence
$E_{2}^{p,q}=\operatorname{H}^{p}(C,\mathbf{R}^{q}f_{*}\mathbb{G}_{m})\implies\operatorname{H}^{p+q}(\mathcal{C},\mathbb{G}_{m}).$
The terms $E_{2}^{p,q}$ for $(p,q)\geq(1,1)$ all vanish, since the sheaves $\mathbf{R}^{q}f_{*}\mathbb{G}_{m}$ are supported on finite sets and hence have vanishing higher cohomology. Additionally, the terms vanish for $(p,0)$ when $p\geq 2$ by Theorem 3.10. Therefore for $k\geq 2$ the only non-zero term in the filtration
$0\subseteq H^{k}F^{k}\subseteq\dots\subseteq H^{0}F^{k}=\operatorname{H}^{k}(\mathcal{C},\mathbb{G}_{m})$
is the term $E_{2}^{0,k}$, and so by Lemma 4.4
$\operatorname{H}^{k}(\mathcal{C},\mathbb{G}_{m})=E_{2}^{0,k}=\operatorname{H}^{0}(C,\mathbf{R}^{k}f_{*}\mathbb{G}_{m})=\bigoplus_{i=1}^{n}\operatorname{H}^{k}(G_{i},\mathbbm{k}^{\times}).$
Note 4.6.
The $p_{i}$’s must be singular points, since the only possible stabilizers of a tame stacky curve at smooth points are cyclic groups.
Proposition 4.7.
The second cohomology of $\mu_{r}$ fits into the following sequence
$\operatorname{H}^{2}(C,\mu_{r})\rightarrow\operatorname{H}^{2}(\mathcal{C},\mu_{r})\rightarrow\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mu_{r})\rightarrow 0.$
That is, a $\mu_{r}$-gerbe over $\mathcal{C}$ is completely determined by a choice of a $\mu_{r}$-gerbe on $C$ and a $\mu_{r}$-gerbe at each non-trivial stabilizer of $\mathcal{C}$.
Proof.
Consider the Leray spectral sequence
$E_{2}^{p,q}=\operatorname{H}^{p}(C,\mathbf{R}^{q}f_{*}\mu_{r})\implies\operatorname{H}^{p+q}(\mathcal{C},\mu_{r}).$
Using Proposition 3.17, the relevant terms in the filtration on $\operatorname{H}^{2}$ are:
$E_{\infty}^{2,0}$ $=E_{3}^{2,0}=\frac{\operatorname{H}^{2}(C,\mu_{r})}{\operatorname{im}\left(E_{2}^{0,1}\rightarrow\operatorname{H}^{2}(C,\mu_{r})\right)}$
$E_{\infty}^{1,1}$ $=0\quad\text{since}\quad E_{2}^{1,1}=\operatorname{H}^{1}(C,\mathbf{R}^{1}f_{*}\mu_{r})=0$
$E_{\infty}^{0,2}$ $=\ker(E_{3}^{0,2}\rightarrow E_{3}^{3,0}=0)$
$=E_{3}^{0,2}=\ker(E_{2}^{0,2}\rightarrow E_{2}^{2,1}=0)$
$=\operatorname{H}^{0}(C,\mathbf{R}^{2}f_{*}\mu_{r}).$
Therefore $\operatorname{H}^{2}(\mathcal{C},\mu_{r})$ surjects onto $\operatorname{H}^{0}(C,\mathbf{R}^{2}f_{*}\mu_{r})=\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mu_{r})$ with kernel equal to a quotient of $\operatorname{H}^{2}(C,\mu_{r})$, as desired. ∎
Corollary 4.8.
A $\mu_{r}$-gerbe on $\mathcal{C}$ is a root gerbe if and only if it is a root gerbe restricted to each $BG_{i}$.
Proof.
Applying the Kummer sequence simultaneously to the three components of the exact sequence in Proposition 4.7 yields
[…]
where the bottom right isomorphism comes from Proposition 4.5. A $\mu_{r}$-gerbe on $\mathcal{C}$ is root if and only if its image in $\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})$ is zero, which happens if and only if, therefore, its image in $\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times})$ is zero. That is, if and only if it is a root gerbe on each $BG_{i}$. ∎
Corollary 4.9.
The $\mu_{r}$-gerbe $\pi:\mathcal{G}\rightarrow\mathcal{C}$ is a root gerbe if and only if $\mathbf{R}^{2}f_{*}\mathbb{G}_{m}\hookrightarrow\mathbf{R}^{2}g_{*}\mathbb{G}_{m}$.
Proof.
By Corollary 4.8 $\mathcal{G}$ is a root gerbe if and only if it is a root gerbe on each fiber. However, by Proposition 3.8 this happens if and only if
$\operatorname{H}^{2}(BG_{i},\mathbb{G}_{m})=\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times})\hookrightarrow\operatorname{H}^{2}(E_{i},\mathbbm{k}^{\times})=\operatorname{H}^{2}(BE_{i},\mathbb{G}_{m})$
for all $i$. By the explicit descriptions of $\mathbf{R}^{2}f_{*}\mathbb{G}_{m}$ and $\mathbf{R}^{2}g_{*}\mathbb{G}_{m}$ given in Lemmas 4.3 and 4.4, this is equivalent to $\mathbf{R}^{2}f_{*}\mathbb{G}_{m}\hookrightarrow\mathbf{R}^{2}g_{*}\mathbb{G}_{m}$. ∎
Lemma 4.10.
There is an injection $\mathbf{R}^{3}f_{*}\mathbb{G}_{m}\hookrightarrow\mathbf{R}^{3}g_{*}\mathbb{G}_{m}$ if and only if $\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})\xrightarrow{\inf}\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$ is injective for all $i$.
Proof.
It is sufficient to check on geometric fibers. The sheaf $\mathbf{R}^{3}f_{*}\mathbb{G}_{m}$ vanishes away from $p_{1},\dots,p_{n}$ by Lemma 4.4, and so it is sufficient to check just at the $p_{i}$. By Corollary 3.15, at each $p_{i}$ we have $(\mathbf{R}^{3}f_{*}\mathbb{G}_{m})_{p_{i}}=\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})$ and $(\mathbf{R}^{3}g_{*}\mathbb{G}_{m})_{p_{i}}=\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$. Therefore $\mathbf{R}^{3}f_{*}\mathbb{G}_{m}\rightarrow\mathbf{R}^{3}g_{*}\mathbb{G}_{m}$ is injective if and only if it is injective on all stalks if and only if the inflation map is injective. ∎
4.3. Leray for $g:\mathcal{G}\rightarrow C$
The terms of the Leray sequence are
$E_{2}^{p,q}=\operatorname{H}^{p}(C,\mathbf{R}^{q}g_{*}\mathbb{G}_{m}).$
Lemma 4.11.
Whenever $p\geq 1$ we have $E_{2}^{p,2}=\operatorname{H}^{p}(C,\mathbf{R}^{2}g_{*}\mathbb{G}_{m})=0$.
Proof.
By Proposition 4.3, $\mathbf{R}^{2}g_{*}\mathbb{G}_{m}$ is supported on finitely many points (the $p_{i}$’s), and so its higher cohomology vanishes. ∎
Lemma 4.12.
Whenever $p\geq 2$, we have $E_{2}^{p,1}=\operatorname{H}^{p}(C,\mathbb{Z}/r)$, and this vanishes for $p\geq 3$.
Proof.
By Lemma 4.1 and Proposition 3.6, we have
$\operatorname{H}^{p}(C,\mathbf{R}^{1}g_{*}\mathbb{G}_{m})=\operatorname{H}^{p}(C,f_{*}\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m})=\operatorname{H}^{p}(C,f_{*}(\mathbb{Z}/r))=\operatorname{H}^{p}(C,\mathbb{Z}/r)$
for all $p\geq 2$. We observe here that by Note 4.2 that in fact this holds for all $p\geq 1$ when $\mathcal{G}\rightarrow\mathcal{C}$ is a root gerbe.
By Theorem 3.10 the Kummer sequence for $C$ at $p\geq 3$ reads
$0=\operatorname{H}^{p-1}(C,\mathbb{G}_{m})\rightarrow\operatorname{H}^{p}(C,\mathbb{Z}/r)\rightarrow\operatorname{H}^{p}(C,\mathbb{G}_{m})=0,$
which shows the vanishing for $p\geq 3$. ∎
Putting all of this together, we get that the $E_{2}$ page of our spectral sequence (with only the relevant terms filled in) looks like
[…]
Proposition 4.13.
The pullback map
$\operatorname{H}^{3}(\mathcal{C},\mathbb{G}_{m})\xrightarrow{\pi^{*}}\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m})$
is injective if and only if the inflation map
$\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})\xrightarrow{\inf}\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$
is injective for all $i$.
Proof.
The associated filtration on $\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m})$ is
$0\subseteq F^{3}H^{3}\subseteq F^{2}H^{3}\subseteq F^{1}H^{3}\subseteq F^{0}H^{3}=\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m}).$
The last term, $F^{0}H^{3}$, surjects onto $E_{\infty}^{0,3}$ with kernel $F^{1}H^{3}$, so we will compute $E_{\infty}^{0,3}$. The relevant differentials are
$d_{2}^{0,3}:E_{2}^{0,3}\rightarrow E_{2}^{2,2},$
$d_{3}^{0,3}:E_{3}^{0,3}\rightarrow E_{3}^{3,1},$
$d_{4}^{0,3}:E_{4}^{0,3}\rightarrow E_{4}^{4,0}.$
Looking at the spectral sequence above this proposition, we see that the only non-zero term appearing in our above differentials is $E_{2}^{0,3}$, and hence
$\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m})\twoheadrightarrow E_{\infty}^{0,3}=E_{2}^{0,3}=\operatorname{H}^{0}(C,\mathbf{R}^{3}g_{*}\mathbb{G}_{m}).$
Using Proposition 4.5 to rewrite $\operatorname{H}^{3}(\mathcal{C},\mathbb{G}_{m})=\operatorname{H}^{0}(C,\mathbf{R}^{3}f_{*}\mathbb{G}_{m})$, we see that the map $\operatorname{H}^{3}(\mathcal{C},\mathbb{G}_{m})\rightarrow\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m})$ is induced by the map $\mathbf{R}^{3}f_{*}\mathbb{G}_{m}\rightarrow\mathbf{R}^{3}g_{*}\mathbb{G}_{m}$. By Lemma 4.10, injectivity of this map is equivalent to injectivity of the inflation map for each $i$.
4.4. Leray for $\pi:\mathcal{G}\rightarrow\mathcal{C}$
We now revisit the Leray spectral sequence for $\pi$ and $\mathbb{G}_{m}$. This will lead to an exact sequence containing $\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})$, and we give multiple conditions which simplify its behavior.
Proposition 4.14.
The sequence of low-degree terms can be extended to
(2) $0\rightarrow\operatorname{H}^{1}(\mathcal{C},\mathbb{G}_{m})\rightarrow\operatorname{H}^{1}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{0}(\mathcal{C},\mathbb{Z}/r)\rightarrow\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})\rightarrow\\ \operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r)\rightarrow\operatorname{H}^{3}(\mathcal{C},\mathbb{G}_{m})\rightarrow\ \operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m}).$
Proof.
This is just the standard exact sequence of low-degree terms, where we add the observations that, since $\mathbf{R}^{2}\pi_{*}\mathbb{G}_{m}=0$,
$\ker\left[\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{0}(\mathcal{C},\mathbf{R}^{2}\pi_{*}\mathbb{G}_{m})\right]=\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})$
and that by Lemma 2.6 we may extend the sequence one additional term to $\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m})$, since $E_{2}^{0,2}=\operatorname{H}^{0}(\mathcal{C},\mathbf{R}^{2}\pi_{*}\mathbb{G}_{m})=0$. ∎
We would like to extract a short exact sequence containing $\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})$ from this. We’ve shown in Lemma 3.8 and Corollary 4.8 that $\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})\rightarrow\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})$ is injective if and only if $\mathcal{G}$ is a root gerbe over $\mathcal{C}$ if and only if $BE_{i}\rightarrow BG_{i}$ is a root gerbe on every fiber, and by Proposition 4.13, we have surjectivity of $\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r)$ if and only if $\operatorname{H}^{3}(\mathcal{C},\mathbb{G}_{m})\rightarrow\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m})$ is injective if and only if the inflation map $\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$ is injective for all $i$. This, combined with Proposition 4.5 to rewrite
$\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})=\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times}),$
yields the following.
Theorem 4.15.
Suppose that $\mathcal{C}$ is a tame separated one-dimensional Deligne-Mumford stack with trivial generic stabilizer of finite type over an algebraically closed field $\mathbbm{k}$. Let $f:\mathcal{C}\rightarrow C$ be the coarse moduli space of $\mathcal{C}$. Suppose that $p_{1},\dots,p_{n}$ are the geometric points in $C$ whose preimage in $\mathcal{C}$ have non-trivial stabilizers $G_{1},\dots,G_{n}$. Let $\pi:\mathcal{G}\rightarrow\mathcal{C}$ be a $\mu_{r}$-gerbe, and denote the residual gerbes in $\mathcal{G}$ over $p_{1},\dots,p_{n}$ as $E_{1},\dots,E_{n}$. Then the following sequence is exact
$\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r).$
Moreover,
  1. (1)
    the left map is injective if and only if $\mathcal{G}$ is a root gerbe over $\mathcal{C}$,
  2. (2)
    $\mathcal{G}$ is a root gerbe over $\mathcal{C}$ if and only if it is a root gerbe on each fiber $BE_{i}\rightarrow BG_{i}$,
  3. (3)
    the right map is surjective if and only if
    $\inf:\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$
    is injective for all $i=1,\dots,n$, and
  4. (4)
    if $\gcd(r,\left|G_{i}\right|)=1$ for all $i=1,\dots,n$, then:
    1. (a)
      $\pi:\mathcal{G}\rightarrow\mathcal{C}$ is automatically a root gerbe,
    2. (b)
      the inflation map is automatically injective, and
    3. (c)
      we have the following splitting
      $\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})=\operatorname{H}^{1}(C,\mathbb{Z}/r)\oplus\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times}).$
Proof.
All that remains to be shown is statement (4). The Kummer sequence on $\mathcal{C}$ reads
$\operatorname{H}^{1}(\mathcal{C},\mathbb{G}_{m})\rightarrow\operatorname{H}^{2}(\mathcal{C},\mu_{r})\rightarrow\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})\xrightarrow{\cdot r}\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m}),$
and so the $\mu_{r}$-gerbe $\mathcal{G}$ on $\mathcal{C}$ (corresponding to the class $[\mathcal{G}]\in\operatorname{H}^{2}(\mathcal{C},\mu_{r})$) is a root gerbe if and only if its image in $\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})$ is trivial. In Proposition 4.5 we have computed that
$\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})=\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times}),$
and all elements in this group are $\prod|G_{i}|$-torsion by Proposition 2.1. But this is coprime to $r$, and all elements of $\operatorname{H}^{2}(\mathcal{C},\mu_{r})$ are $r$-torsion. Therefore the image of any element of $\operatorname{H}^{2}(\mathcal{C},\mu_{r})$ in $\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})$ must vanish, and so $\mathcal{G}$ is a root gerbe. Thus the sequence is injective on the left.
To get surjectivity on the right side of our sequence, it suffices to show that the inflation map $\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$ is injective for all $i$. So consider the Lyndon-Hochschild-Serre spectral sequence for
$0\rightarrow\mu_{r}\rightarrow E_{i}\rightarrow G_{i}\rightarrow 0$
Since $E_{2}^{0,2}=\operatorname{H}^{0}(G_{i},\operatorname{H}^{2}(\mu_{r},\mathbbm{k}^{\times}))=0$ (Proposition 2.2), Lemma 2.6 says that we may extend the sequence of low-degree terms to
$\operatorname{H}^{1}(G_{i},\operatorname{H}^{1}(\mu_{r},\mathbbm{k}^{\times}))=\operatorname{H}^{1}(G_{i},\mathbb{Z}/r)\rightarrow\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$
Note that we cannot conclude that $\operatorname{H}^{1}(G_{i},\mathbb{Z}/r)=\operatorname{Hom}(G_{i},\mathbb{Z}/r)$, since the action may be non-trivial. But we can conclude that it must be annihilated by both $\left|G_{i}\right|$ and $r$ (Proposition 2.1), and so since $\gcd(r,\left|G_{i}\right|)=1$, we see that it is trivial and hence we have injectivity of the inflation map.
Therefore we get an exact sequence, and it splits since both groups are annihilated by coprime integers. Finally, since the stacky structure of $\mathcal{C}$ is coprime to $r$, we see that $\mathbf{R}^{1}f_{*}\mathbb{\mu}_{r}=0$ by Proposition 3.17 and Proposition 2.2, and so we may replace $\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r)$ with $\operatorname{H}^{1}(C,\mathbb{Z}/r)$ (by the sequence of low-degree terms for the Leray spectral sequence for $f:\mathcal{C}\rightarrow C$ and $\mu_{r}$). ∎
Corollary 4.16.
Suppose that $\mathcal{G}$, $\mathcal{C}$, and $C$ are as above. Then if $\mathcal{C}$ is smooth:
  1. (1)
    $\mathcal{G}$ is automatically a root gerbe,
  2. (2)
    the inflation map is automatically injective, and
  3. (3)
    we have
    $\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})=\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r).$
Proof.
By Theorem 3.11 $\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})=0$, which implies that $\mathcal{G}$ is a root gerbe. Now on each stabilizer we have the following exact sequence
$0\rightarrow\mu_{r}\rightarrow E_{i}\rightarrow G_{i}\cong\mathbb{Z}/n_{i}\rightarrow 0$
Since $\mathcal{G}$ is a $\mu_{r}$-banded gerbe the extension is central, which implies that $E_{i}$ is abelian and thus must have a direct summand $\mathbb{Z}/N_{i}$ which surjects onto $\mathbb{Z}/n_{i}$. By Propositions 2.3 and 2.4 we then see that
$\operatorname{H}^{3}(\mathbb{Z}/n_{i},\mathbbm{k}^{\times})\hookrightarrow\operatorname{H}^{3}(\mathbb{Z}/N_{i},\mathbbm{k}^{\times})\hookrightarrow\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times}),$
and so $\operatorname{H}^{3}(\mathcal{C},\mathbb{G}_{m})$ injects into $\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m})$. Therefore
$\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})=\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r).$
The following is just a combination of the two previous statements, but is of enough importance to state on its own. See [Ach24, Proposition 7.10].
Corollary 4.17.
In the case where $\mathcal{C}$ is smooth and $\gcd(r,\left|G_{i}\right|)=1$ for $i=1,\dots,n$, we have
$\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})=\operatorname{H}^{1}(C,\mathbb{Z}/r).$
5. $\operatorname{Br}=\operatorname{Br}^{\prime}=\operatorname{H}^{2}$
Let $\mathcal{X}$ be a separated tame one-dimensional algebraic stack of finite type over a field $\mathbbm{k}$ (here we do not assume that $\mathbbm{k}$ is algebraically closed). Let $\pi:\mathcal{X}\rightarrow X$ be the coarse moduli space of $\mathcal{X}$.
In this section we will prove Theorem 3.4, i.e. that
$\operatorname{Br}(\mathcal{X})=\operatorname{Br}^{\prime}(\mathcal{X})=\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m}).$
Observation 5.1.
Since it is already known that $\operatorname{Br}=\operatorname{Br}^{\prime}$ for a separated tame one-dimensional algebraic stack of finite type over a field ([Mat21, Theorem 2]), it suffices to show that $\operatorname{Br}^{\prime}=\operatorname{H}^{2}$, i.e. that $\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})$ is torsion.
Lemma 5.2.
We have
$\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})\cong\operatorname{H}^{2}(\mathcal{X}_{red},\mathbb{G}_{m}).$
Proof.
Let $I$ be the sheaf of ideals of $\mathcal{X}_{red}$. Since $\mathcal{X}$ is finite type over $\mathbbm{k}$, we may factor $i:\mathcal{X}_{red}\rightarrow\mathcal{X}$ through a sequence of square-zero thickenings and hence may assume that $I$ is itself square-zero.
We then obtain an exact sequence
$1\rightarrow 1+I\rightarrow\mathbb{G}_{m}\rightarrow i_{*}\mathbb{G}_{m}\rightarrow 1$
with associated long exact sequence
$\operatorname{H}^{2}(\mathcal{X},I)\rightarrow\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})\rightarrow\operatorname{H}^{2}(\mathcal{X},i_{*}\mathbb{G}_{m})\rightarrow\operatorname{H}^{3}(\mathcal{X},I).$
We claim that $\operatorname{H}^{2}(\mathcal{X},I)=\operatorname{H}^{3}(\mathcal{X},I)=0$, which would finish the proof as we would then have
$\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})\cong\operatorname{H}^{2}(\mathcal{X},i_{*}\mathbb{G}_{m})\cong\operatorname{H}^{2}(\mathcal{X}_{red},\mathbb{G}_{m}).$
[…]
Therefore we may assume that $\mathcal{X}$ is reduced. Then by [Sta18, Tag 06RC] there is an open $U\subseteq X$ such that $\mathcal{X}_{U}$ is a gerbe. Without loss of generality we may assume that $\mathcal{X}$ is connected, so that for each $x\in U$, the stabilizer $G(x)$ of the preimage of $x$ in $\mathcal{X}$ is a finite group of constant size, say $N$.
The complement of $U$ in $X$ consists of finitely many points, $p_{1},\dots,p_{n}$. Denote by $G_{i}$ the stabilizer of the preimage of $p_{i}$ in $\mathcal{X}$.
Lemma 5.3.
For $p=1,2$, $\mathbf{R}^{p}\pi_{*}\mathbb{G}_{m}$ is torsion.
Proof.
The key is to understand the structure of $(\mathbf{R}^{p}\pi_{*}\mathbb{G}_{m})_{x}$. In Section 3.1 we analyzed this for a tame Deligne-Mumford stack, but we are now in the setting of a tame algebraic stack. Therefore the required methods fall under the purview of [Ach24]. We will show step-by-step how to obtain the result, referring the readers to [Ach24] for proofs. Unless otherwise stated, all following references are to [Ach24]. Corollaries 2.9 and 2.16 together show that
$(\mathbf{R}^{p}\pi_{*}\mathbb{G}_{m})_{x}=\operatorname{H}_{0}^{p}(G(x),\mathbb{G}_{m,R(x)}),$
where $\operatorname{H}_{0}$ denotes Hochschild cohomology, for some finite linearly reductive group $G(x)$ and strictly Henselian local ring $R(x)$. Then Theorem 4.4 (combined with Remark 2.4 to communicate between Hochschild and group cohomology) gives an exact sequence
$0\rightarrow\operatorname{H}^{1}(Q(x),(\mathbbm{k}(x))^{\times})\rightarrow(\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m})_{x}\rightarrow\operatorname{H}_{0}^{1}(\Delta(x),\mathbb{G}_{m})^{Q(x)}\rightarrow$
$\operatorname{H}^{2}(Q(x),(\mathbbm{k}(x))^{\times})\rightarrow(\mathbf{R}^{2}\pi_{*}\mathbb{G}_{m})_{x}\rightarrow 0,$
where
$0\rightarrow\Delta(x)\rightarrow G(x)\rightarrow Q(x)\rightarrow 0$
is the connected-étale sequence of $G(x)$.
Point (2) in the proof of Theorem 4.4 shows that $\operatorname{H}_{0}^{1}(\Delta(x),G_{m})$ is $|\Delta(x)|$-torsion. Further, $\operatorname{H}^{i}(Q(x),(\mathbbm{k}(x))^{\times})$ is $|Q(x)|$-torsion for $i\geq 1$ by Proposition 2.1. This shows that $(\mathbf{R}^{p}\pi_{*}\mathbb{G}_{m})_{x}$ is $|G(x)|$-torsion for $p=1,2$.
If $x\in U$, then $|G(x)|=N$ and hence the stalk is $N$-torsion. If $x$ is not in $U$, then $x=p_{i}$ and then the stalk is $|G_{i}|$-torsion. Therefore the sheaf $\mathbf{R}^{p}\pi_{*}\mathbb{G}_{m}$ is $N\cdot\prod|G_{i}|$-torsion for $p=1,2$. ∎
Theorem (Theorem 3.4).
Let $\mathcal{X}$ be a separated tame one-dimensional algebraic stack of finite type over a field $\mathbbm{k}$ (here we do not assume that $\mathbbm{k}$ is algebraically closed). Then
$\operatorname{Br}(\mathcal{X})=\operatorname{Br}^{\prime}(\mathcal{X})=\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m}).$
Proof.
As noted in Observation 5.1, it suffices to show that $\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})$ is torsion, and as shown in Lemma 5.2 it suffices to assume that $\mathcal{X}$ is reduced. The middle of the sequence of low-degree terms of the Leray spectral sequence for $\pi:\mathcal{X}\rightarrow X$ reads
$\operatorname{H}^{2}(X,\mathbb{G}_{m})\rightarrow\ker(\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})\rightarrow\operatorname{H}^{0}(X,\mathbf{R}^{2}\pi_{*}\mathbb{G}_{m}))\rightarrow\operatorname{H}^{1}(X,\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m}).$
By Lemma 5.3, $\operatorname{H}^{0}(X,\mathbf{R}^{2}\pi_{*}\mathbb{G}_{m})$ and $\operatorname{H}^{1}(X,\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m})$ are torsion. It is classically known ([Gro68, Corollaire 2.2]) that $\operatorname{H}^{2}(X,\mathbb{G}_{m})$ is torsion. Therefore $\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})$ is torsion, and we conclude that
$\operatorname{Br}(\mathcal{X})=\operatorname{Br}^{\prime}(\mathcal{X})=\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m}).$