### v2

# Brauer groups of tame stacky curves and their $\mu_{r}$-gerbes

Abstract.
We fit the Brauer group of a $\mu_{r}$-gerbe over a (possibly arbitrarily singular) tame stacky curve into an exact sequence and give characterizations for when it is short exact and conditions for when it splits. We also give a precise formula for the Brauer group of a $\mu_{r}$-gerbe in the case that the base curve is smooth.
1. Introduction
To date, a significant portion of the current literature on Brauer groups has been dedicated to the following picture: given a specific smooth stacky curve $\mathcal{C}$ and a specific $\mu_{r}$-gerbe $\mathcal{G}$ over $\mathcal{C}$, compute the Brauer group of $\mathcal{G}$. See for instance [AM20], with further progress in [DLP25, Shi19], which studied the Brauer group of $\mathcal{M}_{1,1}$ – a $\mu_{2}$-gerbe over a stacky curve – and laid the foundations for modern work on Brauer groups of stacks. See also [ABJ+24], which studied the Brauer group of $\mathcal{Y}_{0}(2)$, also a $\mu_{2}$-gerbe over a stacky curve. Most work has focused, very successfully, on generalizing the base scheme $S$ for a fixed $\mathcal{C}$ and $\mathcal{G}$.
In this paper, we instead approach this problem from the opposite direction: we fix our base $S$ as an algebraically closed field, and then allow $\mathcal{C}$ to be an arbitrary tame stacky curve and $\mathcal{G}\rightarrow\mathcal{C}$ to be an arbitrary $\mu_{r}$-gerbe, with no smoothness requirements and no requirements on the possible (tame) stabilizers. To this end, our primary results are: complete computations of the Brauer groups of tame stacky curves; complete computations of the Brauer groups of $\mu_{r}$-gerbes over smooth tame stacky curves; and an exact sequence for the Brauer group of a $\mu_{r}$-gerbe over a tame stacky curve along with precise conditions for when it is short-exact and conditions for when it splits.
The main inspiration for, and the fundamental technical basis of, this work is the results in the note [Mei18]. Inspiration was also drawn from [AM20, Lie11, Lop23, Shi21]. While preparing this paper, the author learned of the work [Ach24] which uses a similar approach (leveraging the results of Section 2.1 in conjunction with spectral sequences), although the papers operate on different levels of generality. The paper [Ach24] makes weaker assumptions on the base stack and the étale-ness of stabilizers, offering techniques which work for a wide variety of stacks provided they are locally Brauerless ([Ach24, Definition 4.13]). In the case of abelian stabilizers, this implies that the étale part must be cyclic ([Ach24, Remark 4.10]).
In contrast, this paper works in the more restrictive setting of one-dimensional Deligne-Mumford stacks over algebraically closed fields, but with these stronger assumptions is able to state results for stacks which are highly singular and not locally Brauerless (see Example 3.9 for an interesting example of a stack which is not locally Brauerless). In short, we insist that our stabilizers must be tame and étale, but impose no other conditions. Whereas being locally Brauerless allows for non-étale stabilizers but places strong requirements on the tame étale part. In this way, these two papers work in complementary levels of generality. The research in this paper was conducted independently of [Ach24].
1.1. Main results
Our main result is the following thorough classification of Brauer groups of $\mu_{r}$-gerbes over tame stacky curves. Note that here $\mathcal{C}$ is allowed to be arbitrarily singular.
Theorem (Theorem 4.9).
Suppose that $\mathcal{C}$ is a tame separated one-dimensional Deligne-Mumford stack with trivial generic stabilizer of finite type over an algebraically closed field $\mathbbm{k}$. Let $f:\mathcal{C}\rightarrow C$ be the coarse moduli space of $\mathcal{C}$. Suppose that $p_{1},\dots,p_{n}$ are the geometric points in $C$ whose preimage in $\mathcal{C}$ have non-trivial stabilizers $G_{1},\dots,G_{n}$. Let $\pi:\mathcal{G}\rightarrow\mathcal{C}$ be a $\mu_{r}$-gerbe, and denote the stabilizers in $\mathcal{G}$ over $p_{1},\dots,p_{n}$ as $E_{1},\dots,E_{n}$. Let $G_{i}$ and $E_{i}$ act trivially on $\mathbbm{k}^{\times}$ for each $i$. Then the following sequence is exact, where the left cohomology is group cohomology and the right two are étale cohomology:
$\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r).$
Moreover,
  1. (1)
    the left map is injective if and only if $\mathcal{G}$ is a root gerbe over $\mathcal{C}$,
  2. (2)
    $\mathcal{G}$ is a root gerbe over $\mathcal{C}$ if and only if it is a root gerbe on each fiber $BE_{i}\rightarrow BG_{i}$,
  3. (3)
    the right map is surjective if and only if the inflation map
    $\inf:\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$
    is injective for all $i=1,\dots,n$,
  4. (4)
    if $\gcd(r,\left|G_{i}\right|)=1$ for all $i=1,\dots,n$, then:
    1. (a)
      $\pi:\mathcal{G}\rightarrow\mathcal{C}$ is automatically a root gerbe,
    2. (b)
      the inflation map is automatically injective, and
    3. (c)
      we have the following splitting
      $\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})=\operatorname{H}^{1}(C,\mathbb{Z}/r)\oplus\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times}),$
  5. (5)
    more generally, in the situation that the sequence is right-exact, to show that the sequence splits it suffices to show that on each fiber $BE_{i}\rightarrow BG_{i}$ there is a section
    ${\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times})}$${\operatorname{H}^{2}(E_{i},\mathbbm{k}^{\times}).}$$\scriptstyle{\inf}$
In the case where $\mathcal{C}$ is smooth, this vastly simplifies into the following (the interested reader can find more cases in Section 4.3).
Theorem (Corollary 4.11).
Suppose that $\mathcal{C}$ is a smooth separated tame one-dimensional Deligne-Mumford stack with trivial generic stabilizer of finite type over an algebraically closed field $\mathbbm{k}$. Let $\pi:\mathcal{G}\rightarrow\mathcal{C}$ be a $\mu_{r}$-gerbe. Then
$\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})=\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r).$
We also present a comprehensive computation of the cohomology of $\mathbb{G}_{m}$ on $\mathcal{C}$, with no smoothness requirements.
Theorem (Proposition 3.7).
Let $\mathcal{C}$ be a separated tame one-dimensional Deligne-Mumford stack with trivial generic stabilizer of finite type over an algebraically closed field $\mathbbm{k}$. Let $G_{1},\dots,G_{n}$ be the non-trivial stabilizers of geometric points of $\mathcal{C}$. Then for all $k\geq 2$,
$\operatorname{H}^{k}(\mathcal{C},\mathbb{G}_{m})=\bigoplus_{i=1}^{n}\operatorname{H}^{k}(G_{i},\mathbbm{k}^{\times}),$
where, for even $k$, this is concentrated at the singular points of $\mathcal{C}_{red}$.
Along the way, we also collect multiple characterizations of when a $\mu_{r}$-gerbe over a stacky curve is a root gerbe (see Proposition 2.5 and Corollaries 4.2 and 4.3).
1.2. Structure of the paper
We develop the necessary background on gerbes in Section 2. Section 3 computes the cohomology of $\mathbb{G}_{m}$ on a stacky curve, and Section 4 applies this to prove our main theorem on $\mu_{r}$-gerbes. We collect in Appendix A several results on group cohomology and spectral sequences which are used throughout the paper.
1.3. Conventions
We work over a fixed algebraically closed field $\mathbbm{k}$. All stacks are assumed to be tame, separated, Deligne-Mumford, and of finite type over $\mathbbm{k}$. When $\mathcal{X}$ is one-dimensional, we will refer to $\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})$ as the Brauer group of $\mathcal{X}$. A Deligne-Mumford stack is tame if $\operatorname{char}\mathbbm{k}$ does not divide the order of any geometric stabilizer group, and an algebraic stack is tame if it has finite inertia and all geometric stabilizers are linearly reductive (see [AOV08]).
We take a stacky curve to be a separated one-dimensional Deligne-Mumford stack with trivial generic stabilizer of finite type over a field, though we tend to avoid this terminology (the eagle-eyed reader who has seen the title of this paper will notice that we failed this immediately). Instead, we usually directly state the conditions we are assuming, since there are conflicting definitions of stacky curve in the literature.
We use Lemma 3.1 to assume without loss of generality that all stacks in this paper are reduced.
All groups in this paper act trivially on $\mathbbm{k}^{\times}$ unless otherwise stated.
1.4. Acknowledgments
This project was born out of an upcoming paper with William C. Newman. I am heavily indebted to him for hours of hard work and insightful comments on draft versions of this paper. The results of this paper will be used in this upcoming work to both help understand the structure of ¯$\overline{\mathcal{M}}_{1,1}(B\mu_{n})$ as well as to compute the Picard groups of nodal stacky curves.
I would also like to thank Cherry Ng for helpful conversations on group cohomology, Jackson Morris for fantastic advice on how to compute using spectral sequences, and Minseon Shin for suggesting a simplification of an early version of Proposition 2.11.
2. Basics on gerbes
Definition 2.1.
A stack $\mathcal{X}$ over a site $\mathcal{S}$ is called a gerbe if
  1. (1)
    for every object $T\in\mathcal{S}$, there exists a covering $\{T_{i}\rightarrow T\}$ in $\mathcal{S}$ such that each $\mathcal{X}(T_{i})$ is non-empty, and
  2. (2)
    for objects $x,y\in\mathcal{X}$ over $T\in\mathcal{S}$, there exists a covering $\{T_{i}\rightarrow T\}$ and isomorphisms $x|_{T_{i}}\rightarrow y|_{T_{i}}$ for each $i$.
Definition 2.2.
Let $G$ be an abelian sheaf on $\mathcal{S}$. A stack $\mathcal{X}$ over $\mathcal{S}$ is called a $G$-gerbe (or a banded $G$-gerbe) if it is a gerbe together with the data of isomorphisms $\psi_{x}:G|_{T}\rightarrow\operatorname{Aut}_{T}(x)$ of sheaves for each object $x\in\mathcal{X}(T)$. We require that for each isomorphism $\alpha:x\xrightarrow{\sim}y$ over $T$, the diagram
[…]
commutes.
Loosely speaking, this says that $G$-gerbes are a generalization of the notion of classifying stack, in the sense that a $G$-gerbe is a stack which locally is isomorphic to $BG$, much in the same way that $G$-torsors are locally isomorphic to $G$. It is a foundational fact that $G$-gerbes are classified by $\operatorname{H}^{2}(\mathcal{S},G)$ ([Ols16, Theorem 12.2.8]).
Proposition 2.3.
Let $\pi:\mathcal{X}\rightarrow X$ be a $\mu_{r}$-gerbe over a Deligne-Mumford stack. Then there is a natural isomorphism
$\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m}\cong\operatorname{Hom}(\mu_{r},\mathbb{G}_{m})=\mathbb{Z}/r.$
Proof.
Observe that to each line bundle $L\in\operatorname{Pic}(\mathcal{X})$ we may associate a character $\chi_{L}\in\operatorname{Hom}(\mu_{r},\mathbb{G}_{m})$. For each étale $V\rightarrow X$, we have the exact sequence
$0\rightarrow\operatorname{Pic}(V)\rightarrow\operatorname{Pic}(\mathcal{X}|_{V})\xrightarrow{L\mapsto\chi_{L}}\operatorname{Hom}(\mu_{r},\mathbb{G}_{m}).$
[…]
Our principal gerbes of study will be root gerbes, which were introduced in the literature in [Cad07].
Definition 2.4.
Let $X$ be a Deligne-Mumford stack. Then the data of a line bundle $L$ on $X$ is equivalent to the data of a morphism $X\xrightarrow{[L]}B\mathbb{G}_{m}$. Letting $r$ denote the $r^{\text{th}}$ power map $r:\mathbb{G}_{m}\rightarrow\mathbb{G}_{m}$, the $r^{\text{th}}$ root gerbe, written $X(\sqrt[r]{L})$, is defined as the following fiber product
[…]
A morphism from a scheme $T$ to $X(\sqrt[r]{L})$ is equivalent to the data of a morphism $f:T\rightarrow X$, a line bundle $L^{\prime}$ on $T$, and an isomorphism $L^{\prime\otimes r}\cong f^{*}(L)$.
Proposition 2.5.
Let $X$ be an Deligne-Mumford stack. Then a $\mu_{r}$-gerbe $\pi:\mathcal{X}\rightarrow X$ is a root gerbe if and only if $\operatorname{H}^{2}(X,\mathbb{G}_{m})\rightarrow\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})$ is injective.
Proof.
Consider the Leray spectral sequence
$E_{2}^{p,q}=\operatorname{H}^{p}(X,\mathbf{R}^{q}\pi_{*}\mathbb{G}_{m})\implies\operatorname{H}^{p+q}(\mathcal{X},\mathbb{G}_{m}).$
Since $\pi_{*}\mathbb{G}_{m}=\mathbb{G}_{m}$, the beginning of the associated sequence of low-degree terms is
$0\rightarrow\operatorname{H}^{1}(X,\mathbb{G}_{m})\rightarrow\operatorname{H}^{1}(\mathcal{X},\mathbb{G}_{m})\rightarrow\operatorname{H}^{0}(X,\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m})=$
$\operatorname{H}^{0}(X,\operatorname{Hom}(\mu_{r},\mathbb{G}_{m}))\xrightarrow{d_{2}}\operatorname{H}^{2}(X,\mathbb{G}_{m})\xrightarrow{\pi^{*}}\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m}).$
Therefore the injectivity of $\pi^{*}$ is equivalent to the triviality of $d_{2}$. By [Lop23, Theorem 1.2], $d_{2}$ is just the map which sends $\epsilon:\mu_{r}\rightarrow\mathbb{G}_{m}$ to the image of $[\mathcal{G}]\in\operatorname{H}^{2}(X,\mu_{r})$ under the pushforward map
$\epsilon_{*}:\operatorname{H}^{2}(X,\mu_{r})\rightarrow\operatorname{H}^{2}(X,\mathbb{G}_{m}).$
Setting $\epsilon$ to be the canonical inclusion $\mu_{r}\hookrightarrow\mathbb{G}_{m}$, the Kummer sequence shows that $d_{2}$ is trivial if and only if $[\mathcal{G}]$ is in the image of $\operatorname{H}^{1}(X,\mathbb{G}_{m})$. That is, if and only if $\mathcal{G}$ is a root gerbe. ∎
Proposition 2.6 ([AM20], Proposition 2.5 (iv)).
If $X$ is a regular and noetherian Deligne-Mumford stack, and if $U\subseteq X$ is a dense open subset, then the restriction map $\operatorname{H}^{2}(X,\mathbb{G}_{m})\rightarrow\operatorname{H}^{2}(U,\mathbb{G}_{m})$ is injective.
Theorem 2.7 (Tsen’s Theorem).
Let $X$ be a one-dimensional scheme over an algebraically closed field $\mathbbm{k}$. Then $\operatorname{H}^{i}(X,\mathbb{G}_{m})=0$ for $i\geq 2$.
Theorem 2.8 (Stacky Tsen’s Theorem).
Let $\mathcal{C}$ be a smooth separated one-dimensional Deligne-Mumford stack with trivial generic stabilizer of finite type over an algebraically closed field. Then $\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})=0$.
Proof.
Since there are only finitely many points which have non-trivial stabilizer, this follows from Tsen’s Theorem and Proposition 2.6. ∎
Note 2.9.
Note the contrast between Theorem 2.7, which applies to all curves, and Theorem 2.8, which applies only when $\mathcal{C}$ is smooth. Intuitively, this says that for a stacky curve (which is generically schemey), any non-triviality in $\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})$ must come from stabilizers at singularities, which is precisely what we will see in Proposition 3.7 and Note 3.8. In Example 3.9 we present a minimal example of a stacky curve with non-vanishing $\operatorname{H}^{2}$.
2.1. Some results of Meier
The following results, with small adjustments to fit our context, appear in [ACV03] and [Mei18].
Definition 2.10 ([ATWo20], 2.3.1).
A morphism $f:X\rightarrow Y$ of stacks is called a coarsening if the inertia $\mathcal{I}_{X/Y}$ is finite over $X$ and for any flat $Z\rightarrow Y$ from an algebraic space $Z$ the resulting base change $X\times_{Y}Z\rightarrow Z$ is a coarse moduli space.
Proposition 2.11 ([ACV03], Proposition A.0.1 [Mei18], Proposition 8).
Let $\mathcal{X}$ be a separated tame Deligne-Mumford stack of finite type over $\mathbbm{k}$ with coarsening $f:\mathcal{X}\rightarrow X$. Let $x:\operatorname{Spec}K\rightarrow X$ be a geometric point. Let $G=(\mathcal{I}_{\mathcal{X}/X})_{x}$ be the relative stabilizer. Then for all $n\geq 0$
$(\mathbf{R}^{n}f_{*}\mathbb{G}_{m})_{x}=\operatorname{H}^{n}(G,R^{\times})$
for some strictly Henselian local ring $R$ with residue field $K$. Moreover, the group $G$ acts trivially on $K$.
Proof.
This is originally stated for coarse moduli spaces. We reduce to that case by observing that since $\mathcal{X}\rightarrow X$ is a coarsening we may restrict to any étale neighborhood from a scheme $V\rightarrow X$ and then be in the setting of a coarse moduli space $\mathcal{X}_{V}\rightarrow V$.
This now essentially follows from [ACV03, Proposition A.0.1], which provides an equality
$(\mathbf{R}^{n}f_{*}\mathbb{G}_{m})_{x}=\operatorname{H}^{n}(G,(\mathbb{G}_{m})_{x^{\prime}}),$
where $x^{\prime}$ is the preimage of $x$ in $\mathcal{X}$. All that’s left to do is observe the following: $(\mathbb{G}_{m})_{x^{\prime}}=R^{\times}$ where $R$ is a strictly Henselian local ring; the residue field of $R$ is $K$; and since the relative geometric stabilizer at $X$ is $G$ we must have that $G$ acts trivially on $K$. ∎
Corollary 2.12 ([Mei18], Lemma 9).
Let $\mathcal{X}$ and $G$ be as in Proposition 2.11. Then for all geometric points $x:\operatorname{Spec}K\rightarrow X$ and for all $n\geq 2$,
$(\mathbf{R}^{n}f_{*}\mathbb{G}_{m})_{x}=\operatorname{H}^{n}(G,K^{\times}).$
Note 2.13.
This is originally shown by Meier only for the case $n=2$ and only for the $\ell$-torsion with $\ell\neq\operatorname{char}K$. However, the argument actually works in our stated generality since $\operatorname{H}^{n}(G,-)$ is $|G|$-torsion and $|G|$ is coprime to $\operatorname{char}K$ since $\mathcal{X}$ is tame and Deligne-Mumford. For the sake of completion, we reproduce the argument here.
Proof of Corollary 2.12.
Let $p=\operatorname{char}K$. Using Proposition 2.11, all that’s left to show is that $\operatorname{H}^{n}(G,R^{\times})=\operatorname{H}^{n}(G,K^{\times})$. Consider the exact sequence
$1\rightarrow\ker\rightarrow R^{\times}[1/p]\rightarrow K^{\times}[1/p]\rightarrow 1,$
where if $p=0$ we take adjoining $1/p$ to be trivial. Since $R$ and $K$ are both strictly Henselian, we have that $R^{\times}[1/p]$ and $K^{\times}[1/p]$ are both divisible. Therefore $\ker$ is divisible. Now by [Sta25, Tag 06RR], the torsion of $R^{\times}[1/p]$ maps isomorphically onto the torsion of $K^{\times}[1/p]$, and hence $\ker$ is torsion-free. The structure theorem for divisible groups (see [Fuc70, Theorem 23.1]) then implies that $\ker$ is a $\mathbb{Q}$-vector space and hence has vanishing higher cohomology.
Therefore, taking cohomology, we see that
$\operatorname{H}^{n}(G,R^{\times})=\operatorname{H}^{n}(G,R^{\times}[1/p])=\operatorname{H}^{n}(G,K^{\times}[1/p])=\operatorname{H}^{n}(G,K^{\times})$
for all $n\geq 2$, where the first and last equalities use the fact that $|G|$ is coprime to $p$. ∎
We add here an additional result inspired by the above.
Proposition 2.14.
Let $\mathcal{X}$ and $G$ be as in Proposition 2.11. Then for all geometric points $x:\operatorname{Spec}K\rightarrow X$, all $n\geq 1$, and all $r$ coprime to $\operatorname{char}K$,
$(\mathbf{R}^{n}f_{*}\mu_{r})_{x}=\operatorname{H}^{n}(G,\mu_{r}(K)).$
Proof.
This reduces to showing that if $\mathcal{X}=[\operatorname{Spec}R/G]$, then $\operatorname{H}^{n}(\mathcal{X},\mu_{r})=\operatorname{H}^{n}(G,\mu_{r}(K))$. By the descent spectral sequence we have
$\operatorname{H}^{n}(\mathcal{X},\mu_{r})=\operatorname{H}^{n}(G,\mu_{r}(R))=\operatorname{H}^{n}(G,\mu_{r}(K)),$
where the last equality comes from $R$ being a strictly Henselian local ring. ∎
3. The cohomology of a stacky curve
In this section and the next, we fix the following data:
  • $\mathcal{C}$ is a tame one-dimensional separated Deligne-Mumford stack with trivial generic stabilizer of finite type over an algebraically closed field $\mathbbm{k}$,
  • $\pi:\mathcal{G}\rightarrow\mathcal{C}$ is a $\mu_{r}$-gerbe,
  • $g:\mathcal{G}\rightarrow C$ is the coarse moduli space (note that $C$ is the coarse moduli space of both $\mathcal{G}$ and $\mathcal{C}$),
  • $p_{1},\dots,p_{n}$ are the points of $C$ whose preimage in $\mathcal{C}$ have non-trivial stabilizers $G_{i}$,
  • the preimages of $p_{1},\dots,p_{n}$ in $\mathcal{G}$ have stabilizers $E_{1},\dots,E_{n}$, so that for each $i$ we have the exact sequence of groups
    $0\rightarrow\mu_{r}\rightarrow E_{i}\rightarrow G_{i}\rightarrow 0,$
  • and the trivial action of $G_{i}$ and $E_{i}$ on $\mathbbm{k}^{\times}$ (see Proposition 2.11).
This can be summarized in the diagram
${\mathcal{G}}$${\mathcal{C}}$${C.}$$\scriptstyle{\pi}$$\scriptstyle{g}$$\scriptstyle{f}$
We open by making the following observation.
Lemma 3.1.
Let $\mathcal{X}$ be a separated tame one-dimensional Deligne-Mumford stack of finite type over a (not necessarily algebraically closed field) $\mathbbm{k}$. If $\dim X\leq 1$, then we have
$\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})=\operatorname{H}^{2}(\mathcal{X}_{red},\mathbb{G}_{m}).$
Proof.
Let $I$ be the sheaf of ideals of $\mathcal{X}_{red}$. Since $\mathcal{X}$ is finite type over $\mathbbm{k}$, we may factor $i:\mathcal{X}_{red}\rightarrow\mathcal{X}$ through a sequence of square-zero thickenings and hence may assume that $I$ is itself square-zero.
We then obtain an exact sequence
$1\rightarrow 1+I\rightarrow\mathbb{G}_{m}\rightarrow i_{*}\mathbb{G}_{m}\rightarrow 1$
with associated long exact sequence
$\operatorname{H}^{2}(\mathcal{X},I)\rightarrow\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})\rightarrow\operatorname{H}^{2}(\mathcal{X},i_{*}\mathbb{G}_{m})\rightarrow\operatorname{H}^{3}(\mathcal{X},I).$
We claim that $\operatorname{H}^{2}(\mathcal{X},I)=\operatorname{H}^{3}(\mathcal{X},I)=0$, which would finish the proof as we would then have
$\operatorname{H}^{2}(\mathcal{X},\mathbb{G}_{m})=\operatorname{H}^{2}(\mathcal{X},i_{*}\mathbb{G}_{m})=\operatorname{H}^{2}(\mathcal{X}_{red},\mathbb{G}_{m}).$
[…]
Note 3.2.
We use Lemma 3.1 to without loss of generality treat our stacks as if they were reduced, though we will occasionally directly translate statements into the non-reduced context.
Consider the Leray spectral sequence for the sheaf $\mathbb{G}_{m,\mathcal{G}}$ and the map $\pi:\mathcal{G}\rightarrow\mathcal{C}$:
$E_{2}^{p,q}=\operatorname{H}^{p}(\mathcal{C},\mathbf{R}^{q}\pi_{*}\mathbb{G}_{m,\mathcal{G}})\implies\operatorname{H}^{p+q}(\mathcal{G},\mathbb{G}_{m,\mathcal{G}}).$
Using the fact that $\pi_{*}\mathbb{G}_{m,\mathcal{G}}=\mathbb{G}_{m,\mathcal{C}}$, the associated sequence of low-degree terms is
(1) $0\rightarrow\operatorname{H}^{1}(\mathcal{C},\mathbb{G}_{m})\rightarrow\operatorname{H}^{1}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{0}(\mathcal{C},\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m})\rightarrow\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})\rightarrow\\ \ker\left[\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{0}(\mathcal{C},\mathbf{R}^{2}\pi_{*}\mathbb{G}_{m})\right]\rightarrow\operatorname{H}^{1}(\mathcal{C},\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m})\rightarrow\operatorname{H}^{3}(\mathcal{C},\mathbb{G}_{m}).$
Therefore, before we proceed to our computation of $\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})$, we must first understand the cohomology of $\mathbb{G}_{m}$ on $\mathcal{C}$, as well as $\mathbf{R}^{2}\pi_{*}\mathbb{G}_{m}$. We open our analysis by using the Grothendieck spectral sequence to relate $\mathbf{R}^{*}f_{*}\mathbb{G}_{m}$, $\mathbf{R}^{*}g_{*}\mathbb{G}_{m}$, and $\mathbf{R}^{*}\pi_{*}\mathbb{G}_{m}$.
3.1. Grothendieck spectral sequence
To understand our various higher pushforwards, we consider the Grothendieck spectral sequence for the composition $g=f\circ\pi$ and the sheaf $\mathbb{G}_{m}$:
$E_{2}^{pq}=\left(\mathbf{R}^{p}f_{*}\circ\mathbf{R}^{q}\pi_{*}\right)(\mathbb{G}_{m})\implies\mathbf{R}^{p+q}g_{*}(\mathbb{G}_{m}).$
Accounting for the fact that the pushforward of $\mathbb{G}_{m}$ is $\mathbb{G}_{m}$, the beginning of the sequence of low-degree terms is
[…]
We may extract two short exact sequences from this:
$0\rightarrow\mathbf{R}^{1}f_{*}\mathbb{G}_{m}\rightarrow\mathbf{R}^{1}g_{*}\mathbb{G}_{m}\rightarrow\ker\gamma\rightarrow 0$
and
$0\rightarrow\ker\gamma\rightarrow f_{*}(\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m})\rightarrow\ker\delta\rightarrow 0$
Lemma 3.3.
For all $i\geq 2$,
$\operatorname{H}^{i}(C,\mathbf{R}^{1}g_{*}\mathbb{G}_{m})\cong\operatorname{H}^{i}(C,f_{*}(\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m})).$
The isomorphism holds for $i=1$ if $\mathbf{R}^{2}f_{*}\mathbb{G}_{m}\rightarrow\mathbf{R}^{2}g_{*}\mathbb{G}_{m}$ is injective.
Proof.
Notice that since $\mathcal{C}$ has trivial generic stabilizer, Corollary 2.12 says that for $i\geq 1$ the sheaves $\mathbf{R}^{i}f_{*}\mathbb{G}_{m}$ are supported on a finite set (the points where $\mathcal{C}$ has non-trivial stack structure), and so their higher cohomology vanishes. Hence for $i\geq 1$
$0\rightarrow\operatorname{H}^{i}(C,\mathbf{R}^{1}g_{*}\mathbb{G}_{m})\rightarrow\operatorname{H}^{i}(C,\ker\gamma)\rightarrow 0$
That is,
$\operatorname{H}^{i}(C,\mathbf{R}^{1}g_{*}\mathbb{G}_{m})\cong\operatorname{H}^{i}(C,\ker\gamma).$
But using the second exact sequence from above gives, in degree 0 and 1,
$\operatorname{H}^{0}(C,f_{*}(\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m}))\rightarrow\operatorname{H}^{0}(C,\ker\delta)\rightarrow\operatorname{H}^{1}(C,\ker\gamma)\rightarrow\operatorname{H}^{1}(C,f_{*}(\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m}))\rightarrow 0$
and, in degree above 1,
$0\rightarrow\operatorname{H}^{i}(C,\ker\gamma)\rightarrow\operatorname{H}^{i}(C,f_{*}(\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m}))\rightarrow 0$
This establishes that for all $i\geq 2$ we have
$\operatorname{H}^{i}(C,\mathbf{R}^{1}g_{*}\mathbb{G}_{m})\cong\operatorname{H}^{i}(C,\ker\gamma)\cong\operatorname{H}^{i}(C,f_{*}(\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m}))$
and that the isomorphism holds for $i=1$ if $\operatorname{H}^{0}(C,\ker\delta)=0$, i.e. if $\mathbf{R}^{2}f_{*}\mathbb{G}_{m}\rightarrow\mathbf{R}^{2}g_{*}\mathbb{G}_{m}$ is injective on global sections. In particular this holds if the map of sheaves is injective. ∎
Note 3.4.
We will see in Corollary 4.3 that $\mathbf{R}^{2}f_{*}\mathbb{G}_{m}\rightarrow\mathbf{R}^{2}g_{*}\mathbb{G}_{m}$ is injective precisely when $\mathcal{G}$ is a root gerbe over $\mathcal{C}$, and that in fact the map on sheaves is determined by the map on global sections.
Lemma 3.5.
For all $k\geq 1$,
$\mathbf{R}^{2k}g_{*}\mathbb{G}_{m}=\bigoplus_{i=1}^{n}p_{i_{*}}\operatorname{H}^{2k}(E_{i},\mathbbm{k}^{\times}).$
Proof.
By Corollary 2.12 and Proposition A.2, for all $x\neq p_{1},\dots,p_{n}$ we have $(\mathbf{R}^{2k}g_{*}\mathbb{G}_{m})_{x}=\operatorname{H}^{2k}(\mu_{r},\mathbbm{k}^{\times})=0$. Therefore $\mathbf{R}^{2k}g_{*}\mathbb{G}_{m}$ is supported at the $p_{i}$’s and hence
$\mathbf{R}^{2k}g_{*}\mathbb{G}_{m}=\bigoplus_{i=1}^{n}(\mathbf{R}^{2k}g_{*}\mathbb{G}_{m})_{p_{i}}=\bigoplus_{i=1}^{n}p_{i_{*}}\operatorname{H}^{2k}(E_{i},\mathbbm{k}^{\times}),$
by Corollary 2.12. ∎
3.2. The cohomology of $\mathbb{G}_{m}$ on $\mathcal{C}$
Lemma 3.6.
For all $k\geq 2$,
$\mathbf{R}^{k}f_{*}\mathbb{G}_{m}=\bigoplus_{i=1}^{n}p_{i_{*}}\operatorname{H}^{k}(G_{i},\mathbbm{k}^{\times}).$
Proof.
Let $x$ be a geometric point of $C$ and $G_{x}$ the stabilizer of the preimage of $x$ in $\mathcal{C}$. By Corollary 2.12, $(\mathbf{R}^{k}f_{*}\mathbb{G}_{m})_{x}=\operatorname{H}^{k}(G_{x},\mathbbm{k}^{\times})$, and this vanishes away from $p_{1},\dots,p_{n}$ since $G_{x}=0$ generically. Hence $\mathbf{R}^{k}f_{*}\mathbb{G}_{m}$ is supported at $p_{1},\dots,p_{n}$, and the lemma follows. ∎
Proposition 3.7.
For all $k\geq 2$ we have that
$\operatorname{H}^{k}(\mathcal{C},\mathbb{G}_{m})=\operatorname{H}^{0}(C,\mathbf{R}^{k}f_{*}\mathbb{G}_{m})=\bigoplus_{i=1}^{n}\operatorname{H}^{k}(G_{i},\mathbbm{k}^{\times}),$
where, for even $k$, this is concentrated at the singular points of $\mathcal{C}_{red}$.
Proof.
We use the Leray spectral sequence
$E_{2}^{p,q}=\operatorname{H}^{p}(C,\mathbf{R}^{q}f_{*}\mathbb{G}_{m})\implies\operatorname{H}^{p+q}(\mathcal{C},\mathbb{G}_{m}).$
The terms $E_{2}^{p,q}$ for $(p,q)\geq(1,1)$ all vanish, since the sheaves $\mathbf{R}^{q}f_{*}\mathbb{G}_{m}$ are supported on finite sets (by Corollary 2.12) and hence have vanishing higher cohomology. Additionally, the terms vanish for $(p,0)$ when $p\geq 2$ by Theorem 2.7. Therefore for $k\geq 2$ the only non-zero term in the filtration
$0\subseteq H^{k}F^{k}\subseteq\dots\subseteq H^{0}F^{k}=\operatorname{H}^{k}(\mathcal{C},\mathbb{G}_{m})$
is the term $E_{2}^{0,k}$, and so by Lemma 3.6
$\operatorname{H}^{k}(\mathcal{C},\mathbb{G}_{m})=E_{2}^{0,k}=\operatorname{H}^{0}(C,\mathbf{R}^{k}f_{*}\mathbb{G}_{m})=\bigoplus_{i=1}^{n}\operatorname{H}^{k}(G_{i},\mathbbm{k}^{\times}).$
The last statement follows since the only possible stabilizers at smooth points are cyclic (this is using that $\mathcal{C}$ has trivial generic stabilizer), and cyclic groups have vanishing even cohomology by Proposition A.2. ∎
Note 3.8.
We therefore get a characterization of when restriction to an open $\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})\rightarrow\operatorname{H}^{2}(\mathcal{U},\mathbb{G}_{m})$ fails to be injective: injectivity fails precisely when $\mathcal{C}\setminus\mathcal{U}$ contains a singular point of $\mathcal{C}_{red}$ whose stabilizer has non-trivial second cohomology. This explains where the regularity condition in Proposition 2.6 comes from, and shows that it is truly necessary.
Example 3.9.
Let $G=\mu_{n}\rtimes\mathbb{Z}/2$ and set
$\mathcal{C}=\left[\frac{\mathbbm{k}[x,y]/(xy)}{G}\right],$
where $\mu_{n}$ acts as $t\cdot(x,y)=(tx,t^{a}y)$ such that $a^{2}\equiv 1\pmod{n}$, and where $\mathbb{Z}/2$ acts by interchanging $x$ and $y$. Then $\mathcal{C}$ has a node with a $BG$ at the origin, and its coarse moduli space is $\mathbb{A}^{1}$. A somewhat involved group cohomology computation in an upcoming work with Newman will show that, for certain values of $a$ and $n$, $\operatorname{H}^{2}(G,\mathbbm{k}^{\times})\neq 0$. This gives, in a sense, the simplest possible example of a (generically schemey) stacky curve with non-trivial Brauer group, since a node is the simplest singularity type which supports non-cyclic stabilizers.
4. The Brauer group of a $\mu_{r}$-gerbe
In this section we will complete our computation of the Brauer group of a $\mu_{r}$-gerbe over $\mathcal{C}$. Since such a gerbe corresponds to a class in $\operatorname{H}^{2}(\mathcal{C},\mu_{r})$, we begin with one final cohomology computation on $\mathcal{C}$ to aide in our classification of root gerbes. This will be one of the crucial ingredients in Theorem 4.9.
4.1. Classifying $\mu_{r}$-gerbes over $\mathcal{C}$
Proposition 4.1.
The second cohomology of $\mu_{r}$ fits into the following sequence
$\operatorname{H}^{2}(C,\mu_{r})\rightarrow\operatorname{H}^{2}(\mathcal{C},\mu_{r})\rightarrow\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mu_{r})\rightarrow 0.$
That is, a $\mu_{r}$-gerbe over $\mathcal{C}$ is completely determined by the choice of a $\mu_{r}$-gerbe on $C$ and a $\mu_{r}$-gerbe at each non-trivial stabilizer of $\mathcal{C}$.
Proof.
Consider the Leray spectral sequence
$E_{2}^{p,q}=\operatorname{H}^{p}(C,\mathbf{R}^{q}f_{*}\mu_{r})\implies\operatorname{H}^{p+q}(\mathcal{C},\mu_{r}).$
We use Proposition 2.14 to identify the stalks of $\mathbf{R}^{q}f_{*}\mu_{r}$ with the group cohomology of $\mu_{r}$. Then for $q\geq 1$ the sheaf $\mathbf{R}^{q}f_{*}\mu_{r}$ is supported at finitely many points and hence has vanishing higher cohomology. We then get
$E_{3}^{3,0}=\operatorname{H}^{3}(C,\mu_{r})=0$
by the Kummer sequence,
$E_{2}^{2,1}=\operatorname{H}^{2}(C,\mathbf{R}^{1}f_{*}\mu_{r})=0,\text{ and}$
$E_{2}^{1,1}=\operatorname{H}^{1}(C,\mathbf{R}^{1}f_{*}\mu_{r})=0.$
Then the relevant terms in the filtration on $\operatorname{H}^{2}$ are:
$E_{\infty}^{2,0}$ $=E_{3}^{2,0}=\frac{\operatorname{H}^{2}(C,\mu_{r})}{\operatorname{im}\left(E_{2}^{0,1}\rightarrow\operatorname{H}^{2}(C,\mu_{r})\right)}$
$E_{\infty}^{1,1}$ $=0$
$E_{\infty}^{0,2}$ $=\ker(E_{3}^{0,2}\rightarrow E_{3}^{3,0}=0)$
$=E_{3}^{0,2}=\ker(E_{2}^{0,2}\rightarrow E_{2}^{2,1}=0)$
$=\operatorname{H}^{0}(C,\mathbf{R}^{2}f_{*}\mu_{r}).$
Therefore $\operatorname{H}^{2}(\mathcal{C},\mu_{r})$ surjects onto $\operatorname{H}^{0}(C,\mathbf{R}^{2}f_{*}\mu_{r})=\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mu_{r})$ with kernel equal to a quotient of $\operatorname{H}^{2}(C,\mu_{r})$, as desired. ∎
Corollary 4.2.
A $\mu_{r}$-gerbe on $\mathcal{C}$ is a root gerbe if and only if it is a root gerbe restricted to each $BG_{i}$.
Proof.
Applying the Kummer sequence simultaneously to the three components of the exact sequence in Proposition 4.1 yields
[…]
where the bottom right isomorphism comes from Proposition 3.7. A $\mu_{r}$-gerbe on $\mathcal{C}$ is a root gerbe if and only if its image in $\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})$ is zero, which happens if and only if, therefore, its image in $\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times})$ is zero. That is, if and only if it is a root gerbe on each $BG_{i}$. ∎
Corollary 4.3.
The $\mu_{r}$-gerbe $\pi:\mathcal{G}\rightarrow\mathcal{C}$ is a root gerbe if and only if the natural map $\mathbf{R}^{2}f_{*}\mathbb{G}_{m}\rightarrow\mathbf{R}^{2}g_{*}\mathbb{G}_{m}$ is injective.
Proof.
By Corollary 4.2 $\mathcal{G}$ is a root gerbe if and only if it is a root gerbe on each fiber. However, by Proposition 2.5 this happens if and only if
$\operatorname{H}^{2}(BG_{i},\mathbb{G}_{m})=\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times})\hookrightarrow\operatorname{H}^{2}(E_{i},\mathbbm{k}^{\times})=\operatorname{H}^{2}(BE_{i},\mathbb{G}_{m})$
for all $i$. By the explicit descriptions of $\mathbf{R}^{2}f_{*}\mathbb{G}_{m}$ and $\mathbf{R}^{2}g_{*}\mathbb{G}_{m}$ given in Lemmas 3.5 and 3.6, this is equivalent to $\mathbf{R}^{2}f_{*}\mathbb{G}_{m}\hookrightarrow\mathbf{R}^{2}g_{*}\mathbb{G}_{m}$. ∎
Lemma 4.4.
The natural morphism $\mathbf{R}^{3}f_{*}\mathbb{G}_{m}\rightarrow\mathbf{R}^{3}g_{*}\mathbb{G}_{m}$ is injective if and only if $\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})\xrightarrow{\inf}\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$ is injective for all $i$.
Proof.
It is sufficient to check on geometric fibers. The sheaf $\mathbf{R}^{3}f_{*}\mathbb{G}_{m}$ vanishes away from $p_{1},\dots,p_{n}$ by Lemma 3.6, and so it is sufficient to check just at the $p_{i}$. By Corollary 2.12, at each $p_{i}$ we have $(\mathbf{R}^{3}f_{*}\mathbb{G}_{m})_{p_{i}}=\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})$ and $(\mathbf{R}^{3}g_{*}\mathbb{G}_{m})_{p_{i}}=\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$. Therefore $\mathbf{R}^{3}f_{*}\mathbb{G}_{m}\rightarrow\mathbf{R}^{3}g_{*}\mathbb{G}_{m}$ is injective if and only if it is injective on all stalks if and only if the inflation map is injective. ∎
4.2. Leray for $g:\mathcal{G}\rightarrow C$
The terms of the Leray sequence are
$E_{2}^{p,q}=\operatorname{H}^{p}(C,\mathbf{R}^{q}g_{*}\mathbb{G}_{m}).$
Lemma 4.5.
Whenever $p\geq 1$ we have $E_{2}^{p,2}=\operatorname{H}^{p}(C,\mathbf{R}^{2}g_{*}\mathbb{G}_{m})=0$.
Proof.
By Proposition 3.5, $\mathbf{R}^{2}g_{*}\mathbb{G}_{m}$ is supported on finitely many points (the $p_{i}$’s), and so its higher cohomology vanishes. ∎
Lemma 4.6.
Whenever $p\geq 2$, we have $E_{2}^{p,1}=\operatorname{H}^{p}(C,\mathbb{Z}/r)$, and this vanishes for $p\geq 3$.
Proof.
By Lemma 3.3 and Proposition 2.3, we have
$\operatorname{H}^{p}(C,\mathbf{R}^{1}g_{*}\mathbb{G}_{m})=\operatorname{H}^{p}(C,f_{*}\mathbf{R}^{1}\pi_{*}\mathbb{G}_{m})=\operatorname{H}^{p}(C,f_{*}(\mathbb{Z}/r))=\operatorname{H}^{p}(C,\mathbb{Z}/r)$
for all $p\geq 2$. We observe here that by Lemma 3.3 and Note 3.4 that in fact this holds for all $p\geq 1$ when $\mathcal{G}\rightarrow\mathcal{C}$ is a root gerbe.
By Theorem 2.7 the Kummer sequence for $C$ at $p\geq 3$ reads
$0=\operatorname{H}^{p-1}(C,\mathbb{G}_{m})\rightarrow\operatorname{H}^{p}(C,\mathbb{Z}/r)\rightarrow\operatorname{H}^{p}(C,\mathbb{G}_{m})=0,$
which shows the vanishing for $p\geq 3$. ∎
Putting all of this together, we get that the $E_{2}$ page of our spectral sequence (with only the relevant terms filled in) looks like
[…]
Proposition 4.7.
The pullback map
$\operatorname{H}^{3}(\mathcal{C},\mathbb{G}_{m})\xrightarrow{\pi^{*}}\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m})$
is injective if and only if the inflation map
$\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})\xrightarrow{\inf}\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$
is injective for all $i$.
Proof.
The associated filtration on $\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m})$ is
$0\subseteq F^{3}H^{3}\subseteq F^{2}H^{3}\subseteq F^{1}H^{3}\subseteq F^{0}H^{3}=\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m}).$
The last term, $F^{0}H^{3}$, surjects onto $E_{\infty}^{0,3}$ with kernel $F^{1}H^{3}$, so we will compute $E_{\infty}^{0,3}$. The relevant differentials are
$d_{2}^{0,3}:E_{2}^{0,3}\rightarrow E_{2}^{2,2},$
$d_{3}^{0,3}:E_{3}^{0,3}\rightarrow E_{3}^{3,1},$
$d_{4}^{0,3}:E_{4}^{0,3}\rightarrow E_{4}^{4,0}.$
Looking at the spectral sequence above this proposition, we see that the only non-zero term appearing in our above differentials is $E_{2}^{0,3}$, and hence
$\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m})\twoheadrightarrow E_{\infty}^{0,3}=E_{2}^{0,3}=\operatorname{H}^{0}(C,\mathbf{R}^{3}g_{*}\mathbb{G}_{m}).$
Using Proposition 3.7 to rewrite $\operatorname{H}^{3}(\mathcal{C},\mathbb{G}_{m})=\operatorname{H}^{0}(C,\mathbf{R}^{3}f_{*}\mathbb{G}_{m})$, we see that the map $\operatorname{H}^{3}(\mathcal{C},\mathbb{G}_{m})\rightarrow\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m})$ is induced by the map $\mathbf{R}^{3}f_{*}\mathbb{G}_{m}\rightarrow\mathbf{R}^{3}g_{*}\mathbb{G}_{m}$. By Lemma 4.4, injectivity of this map is equivalent to injectivity of the inflation map for each $i$.
4.3. Leray for $\pi:\mathcal{G}\rightarrow\mathcal{C}$
We now revisit the Leray spectral sequence for $\pi$ and $\mathbb{G}_{m}$. This will lead to an exact sequence containing $\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})$, and we give multiple conditions which simplify its behavior.
Proposition 4.8.
The sequence of low-degree terms can be extended to
(2) $0\rightarrow\operatorname{H}^{1}(\mathcal{C},\mathbb{G}_{m})\rightarrow\operatorname{H}^{1}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{0}(\mathcal{C},\mathbb{Z}/r)\rightarrow\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})\rightarrow\\ \operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r)\rightarrow\operatorname{H}^{3}(\mathcal{C},\mathbb{G}_{m})\rightarrow\ \operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m}).$
Proof.
This is just the standard exact sequence of low-degree terms (Sequence 1), where we add the following observations.
First, we have $\mathbf{R}^{2}\pi_{*}\mathbb{G}_{m}=0$ since for all geometric points $x$ of $\mathcal{C}$ we have $(\mathbf{R}^{2}\pi_{*}\mathbb{G}_{m})_{x}=\operatorname{H}^{2}(\mu_{r},\mathbbm{k}^{\times})=0$ by Corollary 2.12 and Proposition A.2. Therefore
$\ker\left[\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{0}(\mathcal{C},\mathbf{R}^{2}\pi_{*}\mathbb{G}_{m})\right]=\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m}).$
Additionally, by Lemma A.6 we may extend the sequence one additional term to $\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m})$, since $E_{2}^{0,2}=\operatorname{H}^{0}(\mathcal{C},\mathbf{R}^{2}\pi_{*}\mathbb{G}_{m})=0$. ∎
We would like to extract a short exact sequence containing $\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})$ from this. We’ve shown in Lemma 2.5 and Corollary 4.2 that $\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})\rightarrow\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})$ is injective if and only if $\mathcal{G}$ is a root gerbe over $\mathcal{C}$ if and only if $BE_{i}\rightarrow BG_{i}$ is a root gerbe on every fiber. By Proposition 4.7, we have surjectivity of $\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r)$ if and only if $\operatorname{H}^{3}(\mathcal{C},\mathbb{G}_{m})\rightarrow\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m})$ is injective if and only if the inflation map $\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$ is injective for all $i$. This, combined with Proposition 3.7 to rewrite
$\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})=\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times}),$
yields the first three statements of the following.
Theorem 4.9.
Suppose that $\mathcal{C}$ is a tame separated one-dimensional Deligne-Mumford stack with trivial generic stabilizer of finite type over an algebraically closed field $\mathbbm{k}$. Let $f:\mathcal{C}\rightarrow C$ be the coarse moduli space of $\mathcal{C}$. Suppose that $p_{1},\dots,p_{n}$ are the geometric points in $C$ whose preimage in $\mathcal{C}$ have non-trivial stabilizers $G_{1},\dots,G_{n}$. Let $\pi:\mathcal{G}\rightarrow\mathcal{C}$ be a $\mu_{r}$-gerbe, and denote the stabilizers in $\mathcal{G}$ over $p_{1},\dots,p_{n}$ as $E_{1},\dots,E_{n}$. Let $G_{i}$ and $E_{i}$ act trivially on $\mathbbm{k}^{\times}$ for each $i$. Then the following sequence is exact
$\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})\rightarrow\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r).$
Moreover,
  1. (1)
    the left map is injective if and only if $\mathcal{G}$ is a root gerbe over $\mathcal{C}$,
  2. (2)
    $\mathcal{G}$ is a root gerbe over $\mathcal{C}$ if and only if it is a root gerbe on each fiber $BE_{i}\rightarrow BG_{i}$,
  3. (3)
    the right map is surjective if and only if the inflation map
    $\inf:\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$
    is injective for all $i=1,\dots,n$,
  4. (4)
    if $\gcd(r,\left|G_{i}\right|)=1$ for all $i=1,\dots,n$, then:
    1. (a)
      $\pi:\mathcal{G}\rightarrow\mathcal{C}$ is automatically a root gerbe,
    2. (b)
      the inflation map is automatically injective, and
    3. (c)
      we have the following splitting
      $\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})=\operatorname{H}^{1}(C,\mathbb{Z}/r)\oplus\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times}),$
  5. (5)
    more generally, in the situation that the sequence is right-exact, to show that the sequence splits it suffices to show that on each fiber $BE_{i}\rightarrow BG_{i}$ there is a section
    ${\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times})}$${\operatorname{H}^{2}(E_{i},\mathbbm{k}^{\times}).}$$\scriptstyle{\inf}$
Proof.
All that remains to be shown are statements (4) and (5).
For (4), The Kummer sequence on $\mathcal{C}$ reads
$\operatorname{H}^{1}(\mathcal{C},\mathbb{G}_{m})\rightarrow\operatorname{H}^{2}(\mathcal{C},\mu_{r})\rightarrow\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})\xrightarrow{\cdot r}\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m}),$
and so the $\mu_{r}$-gerbe $\mathcal{G}$ on $\mathcal{C}$ (corresponding to the class $[\mathcal{G}]\in\operatorname{H}^{2}(\mathcal{C},\mu_{r})$) is a root gerbe if and only if its image in $\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})$ is trivial. In Proposition 3.7 we have computed that
$\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})=\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times}),$
and all elements in this group are $\prod|G_{i}|$-torsion by Proposition A.1. But this is coprime to $r$, and all elements of $\operatorname{H}^{2}(\mathcal{C},\mu_{r})$ are $r$-torsion. Therefore the image of any element of $\operatorname{H}^{2}(\mathcal{C},\mu_{r})$ in $\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})$ must vanish, and so $\mathcal{G}$ is a root gerbe. Thus the sequence is injective on the left.
To get surjectivity on the right side of our sequence, it suffices to show that the inflation map $\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$ is injective for all $i$. So consider the Lyndon-Hochschild-Serre spectral sequence for
$0\rightarrow\mu_{r}\rightarrow E_{i}\rightarrow G_{i}\rightarrow 0$
Since $E_{2}^{0,2}=\operatorname{H}^{0}(G_{i},\operatorname{H}^{2}(\mu_{r},\mathbbm{k}^{\times}))=0$ (by Proposition A.2), Lemma A.6 says that we may extend the sequence of low-degree terms to
$\operatorname{H}^{1}(G_{i},\operatorname{H}^{1}(\mu_{r},\mathbbm{k}^{\times}))=\operatorname{H}^{1}(G_{i},\mathbb{Z}/r)\rightarrow\operatorname{H}^{3}(G_{i},\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times})$
Note that we cannot conclude that $\operatorname{H}^{1}(G_{i},\mathbb{Z}/r)=\operatorname{Hom}(G_{i},\mathbb{Z}/r)$, since the action may be non-trivial. But we can conclude that it must be annihilated by both $\left|G_{i}\right|$ and $r$ (Proposition A.1), and so since $\gcd(r,\left|G_{i}\right|)=1$, we see that it is trivial and hence we have injectivity of the inflation map.
Therefore we get an exact sequence, and it splits since both groups are annihilated by coprime integers. Finally, since the stacky structure of $\mathcal{C}$ is coprime to $r$, we see that $\mathbf{R}^{1}f_{*}\mathbb{\mu}_{r}=0$ by Proposition 2.14 and Proposition A.2, and so we may replace $\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r)$ with $\operatorname{H}^{1}(C,\mathbb{Z}/r)$ (by the sequence of low-degree terms for the Leray spectral sequence for $f:\mathcal{C}\rightarrow C$ and $\mu_{r}$).
To see (5), we assume that the sequence is right-exact and that we have sections
${\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times})}$${\operatorname{H}^{2}(E_{i},\mathbbm{k}^{\times}).}$$\scriptstyle{\pi^{*}}$
on each fiber, and we seek to show the sequence splits. But we simply observe there is an induced splitting from the following diagram
[…]
Remark 4.10.
We give here a quick description of the left map.
Let $(\alpha_{i})\in\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times})$. Since
$\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})=\bigoplus_{i=1}^{n}\operatorname{H}^{2}(G_{i},\mathbbm{k}^{\times})$
we know that there is a unique $\mathbb{G}_{m}$-gerbe over $\mathcal{C}$, which we will call $\mathcal{H}$, whose restriction to each $BG_{i}$ is the $\mathbb{G}_{m}$-gerbe $\alpha_{i}$. Then the image of $(\alpha_{i})$ in $\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})$ is just $\pi^{*}([\mathcal{H}])$, i.e. $[\mathcal{G}\times_{\mathcal{C}}\mathcal{H}]$. This is a $\mathbb{G}_{m}$-gerbe which is trivial away from the $BE_{i}$ and whose restriction to each $BE_{i}$ is given by $\inf\alpha_{i}\in\operatorname{H}^{2}(E_{i},\mathbbm{k}^{\times})$.
Corollary 4.11.
Suppose that $\mathcal{G}$, $\mathcal{C}$, and $C$ are as above. Then if $\mathcal{C}$ is smooth:
  1. (1)
    $\mathcal{G}$ is automatically a root gerbe,
  2. (2)
    the inflation map is automatically injective, and
  3. (3)
    we have
    $\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})=\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r).$
Proof.
By Theorem 2.8 $\operatorname{H}^{2}(\mathcal{C},\mathbb{G}_{m})=0$, which implies that $\mathcal{G}$ is a root gerbe. Now on each stabilizer we have the following exact sequence
$0\rightarrow\mu_{r}\rightarrow E_{i}\rightarrow G_{i}\cong\mathbb{Z}/n_{i}\rightarrow 0$
Since $\mathcal{G}$ is a $\mu_{r}$-banded gerbe the extension is central. Therefore $E_{i}$ is abelian – as it’s a central extension of a cyclic group by a cyclic group – and thus must have a direct summand $\mathbb{Z}/N_{i}$ which surjects onto $\mathbb{Z}/n_{i}$. By Propositions A.3 and A.4 we then see that
$\operatorname{H}^{3}(\mathbb{Z}/n_{i},\mathbbm{k}^{\times})\hookrightarrow\operatorname{H}^{3}(\mathbb{Z}/N_{i},\mathbbm{k}^{\times})\hookrightarrow\operatorname{H}^{3}(E_{i},\mathbbm{k}^{\times}),$
and so $\operatorname{H}^{3}(\mathcal{C},\mathbb{G}_{m})$ injects into $\operatorname{H}^{3}(\mathcal{G},\mathbb{G}_{m})$. Therefore
$\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})=\operatorname{H}^{1}(\mathcal{C},\mathbb{Z}/r).$
The following is just a combination of the two previous statements, but is of enough importance to state on its own. See [Ach24, Proposition 7.10].
Corollary 4.12.
In the case where $\mathcal{C}$ is smooth and $\gcd(r,\left|G_{i}\right|)=1$ for $i=1,\dots,n$, we have
$\operatorname{H}^{2}(\mathcal{G},\mathbb{G}_{m})=\operatorname{H}^{1}(C,\mathbb{Z}/r).$
Appendix A Assorted background
This paper relies very heavily on the careful application of both spectral sequences and group cohomology. Therefore we consider it appropriate to give a brief explanation of the properties and results invoked above.
A.1. Group cohomology
Most of this can be found in a standard reference, such as [DF04, Chapter 17] or [Wei94, Chapter 6].
For a group $G$ and an abelian group $A$ on which $G$ acts additively (called a $G$-module), define the invariant subgroup
$A^{G}=\{a\in A:ga=a,\ \forall g\in G\}.$
Then the cohomology of $G$ with coefficients in $A$, $\operatorname{H}^{n}(G,A)$, is the right derived functor $\mathbf{R}^{n}(-^{G})(A)$.
Proposition A.1.
Let $G$ be a group and $A$ a $G$-module. If $G$ is $m$-torsion, then so is $\operatorname{H}^{i}(G,A)$ for all $i\geq 1$. If $A$ is $n$-torsion, then so is $\operatorname{H}^{i}(G,A)$ for all $i\geq 0$.
Proposition A.2.
Let $\mathbbm{k}$ be an algebraically closed field. Let $\mu_{r}$ act trivially on $\mathbbm{k}^{\times}$. Then
$\operatorname{H}^{2k-1}(\mu_{r},\mathbbm{k}^{\times})=\operatorname{Hom}(\mu_{r},\mathbbm{k}^{\times})=\mathbb{Z}/r$
and
$\operatorname{H}^{2k}(\mu_{r},\mathbbm{k}^{\times})=0$
for $k\geq 1$.
Proposition A.3.
Let $\mathbb{Z}/n$ and $\mathbb{Z}/mn$ act trivially on $\mathbbm{k}^{\times}$. Then
$\inf:\operatorname{H}^{3}(\mathbb{Z}/n,\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{3}(\mathbb{Z}/mn,\mathbbm{k}^{\times})$
is given by the canonical injection $1\mapsto m$.
Proof.
More generally, the map
$\operatorname{H}^{k}(\mathbb{Z}/n,\mathbbm{k}^{\times})\rightarrow\operatorname{H}^{k}(\mathbb{Z}/mn,\mathbbm{k}^{\times})$
is given by $1\mapsto m^{\left\lfloor{\frac{k}{2}}\right\rfloor}$ [Ser79, XIII Ex. 1.1], which for $k=3$ is $1\mapsto m$. ∎
Proposition A.4.
If $\mathbb{Z}/a$ and $\mathbb{Z}/b$ both act trivially on $\mathbbm{k}^{\times}$, then
$\operatorname{H}^{k}(\mathbb{Z}/a,\mathbbm{k}^{\times})\hookrightarrow\operatorname{H}^{k}(\mathbb{Z}/a\oplus\mathbb{Z}/b,\mathbbm{k}^{\times})$
for all $k\geq 1$.
A.2. Spectral Sequences
A standard reference for this section is [Wei94, Chapter 5].
Definition A.5.
A spectral sequence consists of the following data
  1. (1)
    sheets $E_{r}$ for $r\geq r_{0}$,
  2. (2)
    for all integers $p,q$, objects $E_{r}^{p,q}$, and
  3. (3)
    maps $d_{r}:E_{r}^{p,q}\rightarrow E_{r}^{p+r,q-(r-1)}$ such that $d_{r}\circ d_{r}=0$ and $E_{r+1}^{p,q}$ is the homology of $E_{r}$ at $(p,q)$.
In practice, many spectral sequences (including the ones used in this paper), begin at the $E_{2}$ sheet. One then generally uses the $E_{2}$ sheet to calculate the subsequent sheets.
In this section we will assume that all spectral sequences are first quadrant, meaning the only non-zero terms have $p,q\geq 0$. Note that any first quadrant spectral sequence must eventually stabilize in the sense that there is some $N$ such that $E_{r}^{p,q}=E_{N}^{p,q}$ for all $r\geq N$. We call this stable value $E_{\infty}^{p,q}$.
We say that a first quadrant spectral sequence converges to $\operatorname{H}^{*}$ for some collection of objects $(\operatorname{H}^{n})$ if for each $n$ there is a filtration
$0\subseteq F^{n}\operatorname{H}^{n}\subseteq F^{n-1}\operatorname{H}^{n}\subseteq\dots\subseteq F^{0}\operatorname{H}^{n}=\operatorname{H}^{n}$
such that
$E_{\infty}^{p,q}\cong F^{p}\operatorname{H}^{p+q}/F^{p+1}\operatorname{H}^{p+q}.$
When such a filtration exists, we denote this by
$E_{r}^{p,q}\implies\operatorname{H}^{p+q}.$
Convergent first quadrant spectral sequences always have a sequence of low-degree terms.
Lemma A.6 (Sequence of low-degree terms).
If
$E_{2}^{p,q}\implies\operatorname{H}^{p+q}(X)$
is a first quadrant spectral sequence, then there is an exact sequence
$0\rightarrow E_{2}^{1,0}\rightarrow\operatorname{H}^{1}(X)\rightarrow E_{2}^{0,1}\rightarrow E_{2}^{2,0}\rightarrow\ker\left[\operatorname{H}^{2}(X)\rightarrow E_{2}^{0,2}\right]\rightarrow E_{2}^{1,1}\rightarrow E_{2}^{3,0}.$
This can be extended an additional term to $\operatorname{H}^{3}(X)$ if $E_{2}^{0,2}=0$.
Proof.
This is standard, but we will give an explanation of the extension to $\operatorname{H}^{3}(X)$. The filtration on $\operatorname{H}^{3}(X)$ associated to the spectral sequence is
$0\subseteq F^{3}\operatorname{H}^{3}\subseteq F^{2}\operatorname{H}^{3}\subseteq F^{1}\operatorname{H}^{3}\subseteq F^{0}\operatorname{H}^{3}=\operatorname{H}^{3}(X),$
where $F^{i}\operatorname{H}^{3}/F^{i+1}\operatorname{H}^{3}\cong E_{\infty}^{i,3-i}$. Therefore there is a canonical injection $E_{\infty}^{3,0}\hookrightarrow\operatorname{H}^{3}(X)$. Exactness of our desired extension says that we need a morphism $E_{2}^{3,0}\rightarrow\operatorname{H}^{3}(X)$ whose kernel is precisely the image of $E_{2}^{1,1}$.
Note that
$E_{3}^{3,0}=\frac{\ker(E_{2}^{3,0}\rightarrow E_{2}^{5,-1}=0)}{\operatorname{im}(E_{2}^{1,1}\rightarrow E_{2}^{3,0})}$
and so we have exactness of the extension as long as $E_{3}^{3,0}=E_{\infty}^{3,0}$. The obstruction to this happening is whether or not $E_{4}^{3,0}=E_{3}^{3,0}$. Since
$E_{4}^{3,0}=\frac{\ker(E_{3}^{3,0}\rightarrow E_{3}^{6,-2}=0)}{\operatorname{im}(E_{3}^{0,2}\rightarrow E_{3}^{3,0})},$
we see that $E_{4}^{3,0}=E_{3}^{3,0}$ if and only if the image of $E_{3}^{0,2}$ is trivial in $E_{3}^{3,0}$. In particular, we get equality and hence a further extension of the sequence if $E_{2}^{0,2}=0$. ∎
We will make extensive use of the Grothendieck spectral sequence, the Leray spectral sequence, and the Lyndon-Hochschild-Serre spectral sequence. Our statement of the Grothendieck spectral sequence is only for the context in which we will use it (pushforward of abelian sheaves on stacks), though it applies in much larger generality.
Definition A.7.
For morphisms $f:\mathcal{X}\rightarrow\mathcal{Y}$ and $g:\mathcal{Y}\rightarrow\mathcal{Z}$ of stacks and an abelian sheaf $F$ on $\mathcal{X}$, the Grothendieck spectral sequence is
$E_{2}^{p,q}=\mathbf{R}^{p}g_{*}(\mathbf{R}^{q}f_{*}F)\implies\mathbf{R}^{p+q}(g\circ f)_{*}F.$
Definition A.8.
For a morphism $f:\mathcal{X}\rightarrow\mathcal{Y}$ of stacks and an abelian sheaf $F$ on $\mathcal{X}$, the Leray spectral sequence is
$E_{2}^{p,q}=\operatorname{H}^{p}(\mathcal{Y},\mathbf{R}^{q}f_{*}F)\implies\operatorname{H}^{p+q}(\mathcal{X},F).$
Definition A.9.
For a group $G$ with normal subgroup $N$ and $G$-module $A$, the Lyndon-Hochschild-Serre spectral sequence is
$E_{2}^{p,q}=\operatorname{H}^{p}(G/N,\operatorname{H}^{q}(N,A))\implies\operatorname{H}^{p+q}(G,A).$